### v1

# Covariate balancing estimation and model selection for difference-in-differences approach

Abstract
In causal inference, remarkable progress has been made in difference-in-differences (DID) approaches to estimate the average effect of treatment on the treated (ATT). Of these, the semiparametric DID (SDID) approach incorporates a propensity score analysis into the DID setup. Supposing that the ATT is a function of covariates, we estimate it by weighting the inverse of the propensity score. As one method to make the estimation robust to the propensity score modeling, we incorporate covariate balancing. Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator has doubly robustness. In addition to the estimation, model selection is also addressed. In practice, covariate selection is an essential task in statistical analysis, but even in the basic setting of the SDID approach, there are no reasonable information criteria. Therefore, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation. As a result, we show that a penalty term is derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria. Numerical experiments show that the proposed method estimates the ATT robustly compared to the method using propensity scores given by the maximum likelihood estimation (MLE), and that the proposed criterion clearly reduces the risk targeted in the SDID approach compared to the intuitive generalization of the existing information criterion. In addition, real data analysis confirms that there is a large difference between the results of the proposed method and the existing method.
Keywords: Covariate balancing, Difference-in-differences, Information criterion, Model selection, Propensity score, Semiparametric approach, Statistical asymptotic theory
1 Introduction
The difference-in-differences (DID) approach is one of the most common methods used in fields such as economics and epidemiology to assess the effect of a particular intervention or treatment. In the basic setting, there are two groups: a treatment group receiving a specific treatment and a control group receiving no treatment or other treatment. The pre and post-treatment outcomes of each group are measured and the differences in the outcomes are compared between the groups to estimate the average treatment effect on the treated (ATT). Under the parallel trend assumption, simple estimation is valid. More recently, Abadie (2005) dealt with inference under more naturalized assumptions using information from the covariates, specifically the parallel trend assumption conditioned on the covariates. Then, as a method that does not necessarily require modeling of the covariates to the outcome, they propose the semiparametric DID (SDID) method, which estimates ATT by weighting the inverse of the propensity score.
With or without conditionality, the parallel trend assumption is a key assumption in the DID. This is the assumption that the mean outcomes of a treatment group and a control group would have been similar over time if there had been no treatment. On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g., Robins et al. 1994 and Li et al. 2018). This is the assumption that the potential outcome variable and the treatment assignment variable, which represents the received treatment, become independent conditioned on the covariates. Although neither assumption is verifiable, the validity of the parallel trend assumption can be examined by visual inspection when we have observations over many time points. And the parallel trend assumption naturally holds if it can be assumed that there is no change over time in the expected value of the outcome without the treatment. Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such desirable conditions is the advantage that the DID gained in exchange for collecting pre- and post-treatment outcomes.
The methodological development of the DID has been remarkable in recent years because of such advantages and the increasing amount of available observational study data (see Roth et al. 2023 for a comprehensive review of the DID in recent years). However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified, but if the model is incorrect, the resulting estimator has a bias. As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome, and proposes the estimator for the ATT that combines the results of each estimator. Their proposed estimator is consistent if either a propensity score or outcome regression working model is correctly specified, that is, the estimators have doubly robustness.
In this article, as with Abadie (2005), we assume that ATT is a function of covariates and address the semiparametric estimation problem for that function. We assume two regression models as above and give double robustness to the estimation, but unlike the Sant’Anna and Zhao (2020), we incorporate covariate balancing (Imai and Ratkovic 2014), a method that uses propensity scores such that a certain moment condition is satisfied. This method actually estimates the model for the treatment assignment, while assuming, but not estimating, the model for the outcome. Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either one of the regression models is correctly specified. One point to note is that, in order to satisfy this favorable asymptotic property in DID, the moment condition is constructed in a different way from the usual one. Incidentally, incorporating covariate balancing leads to the estimation of causal effects via weights that always balance the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014). We call the proposed methodology the covariate balancing for DID (CBD) method.
Obviously, model selection is an essential task in the DID as well, in a series of data analysis tasks. For example, in estimating the effect of union membership on wages, Houngbedji (2016) treats a linear model with workers’ demographics for causal effects, and explores factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value. In other words, the issue of covariate selection is addressed. However, for the information criterion that is the basic tool for model selection, there are no reasonable ones, even for the basic setting of the SDID. Thus, we derive the information criterion for the basic SDID and CBD methods as an asymptotically unbiased estimator of the risk based on the loss function used in the SDID estimation. Note that traditional methods such as GIC (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which is a random variable. Since the SDID method places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment condition when deriving the information criterion. This causes some difficulties. As a result, we derive penalty terms that are significantly different from twice the number of parameters, which often appear in AIC-type information criteria including GIC.
In summary, the aims of this article are to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID method, and to derive model selection criteria for the SDID method, for which no reasonable model selection criteria exist yet. The organization of this article is as follows. First, in section 2, the model and estimation target are specified, and the SDID method proposed in Abadie (2005) is introduced. In section 3, we propose the CBD method and give a theorem showing its doubly robustness. Numerical experiments show that the estimation is robust against model misspecification. In section 4, when the propensity score is known, we derive the model selection criterion as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID method. Then, we also give a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid. As a comparison, we derive the model selection criterion the QICw, which naturally extends the idea of Platt et al. (2013) to the SDID method. Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of estimating the risk, and also verify its performance compared to the QICw from the viewpoint of risk minimization. Section 5 summarizes the results for the CBD method in the same manner as Section 4. In other words, model selection criteria are derived, theorems are given to support their validity, and their performance is evaluated through numerical experiments. Section 6 reports the results of estimating causal effects and model selection for the real data in LaLonde (1986). Finally, Section 7 discusses the future development of the proposed method.
2 Preparation
2.1 Model and Assumptions
Suppose that there are two groups, a treatment group and a control group, and that outcomes are collected before and after the possible treatment time point, respectively. We treat the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome of the treatment group, $y^{[0]}(t)$ the outcome of the control group, $t=0$ the outcome before treatment, and $t=1$ the outcome after treatment. In practice, only one of the outcomes of the treatment or control group can be obtained. Preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if assigned to the control group, the observed outcome collected at each time point can be expressed as
$y(t)=d^{[1]}y^{[1]}(t)+d^{[0]}y^{[0]}(t).$
Letting the change in the outcome variable at two time points in the treatment group be $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group be $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change can be expressed as
$\Delta=d^{[1]}\Delta^{[1]}+d^{[0]}\Delta^{[0]}.$
Here, for the ATT conditioned on the covariate $\bm{x}\ (\in\mathbb{R}^{p})$,
$a(\bm{x})\equiv{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1],$ (1)
we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, we suppose the working model to be a linear model $\bm{x}^{{\rm T}}\bm{\theta}$. Then, we estimate $\bm{\theta}$ and give the ATT by taking the expectation with respect to $\bm{x}$ in the treatment group.
Now, we give the optimal value of $\bm{\theta}$ as
$\bm{\theta}^{*}\equiv\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[\{{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}\mid d^{[1]}=1].$ (2)
Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated by the SDID method using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and thus the ATT can be consistently estimated. The assumptions are as follows.
Assumption 1 (Conditional parallel trend).
${\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[0]}=1],$
that is, for the expected value conditioned on $\bm{x}$ of the change in outcome over time without the treatment, the values for the treatment and control groups are the same.
Assumption 2 (Common shocks).
Events occurring between the time points of outcome collection before and after the treatment have the same impact on outcomes in both the treatment group and the control group.
Assumption 3 (Positivity).
$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1,$
that is, the propensity score is not 00 or $1$.
Abadie (2005) showed that using the weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$, then
${\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]$ $={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]$
$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]$
$={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]$
$=a(\bm{x})$ (3)
holds. Assumption 1 is used in the second equation. This equation means that by multiplying weights based on the propensity scores to the observed values of change, the conditional ATT can be estimated via DID. Then, from (2) and (3), it is shown that the optimal solution of $\bm{\theta}$ can be expressed as
$\bm{\theta}^{*}=\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[e^{[1]}(\bm{x})\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}].$ (4)
From this equation,
${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$ $={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$=\bm{0}$ (5)
is satisfied, and therefore, the equation
${\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$ (6)
is obtained.
2.2 Semiparametric estimation
Given an i.i.d. sample of size $n$, the $i$-th sample is denoted with the subscript $i$. Abadie (2005) proposed
$\hat{\bm{\theta}}=\Bigg{\{}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\Bigg{\}}^{-1}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ (7)
as a natural estimator of $\bm{\theta}^{*}$ by solving (4). When the weights based on the propensity scores are multiplied to the differences of the observed values, a term corresponding to the difference in difference appears, thus this method is called the Semiparametric DID (SDID) estimator. The following regular conditions are assumed.
Assumption 4.
$(\mathrm{i})$ $\bm{\theta}^{*}$ is interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$. $(\mathrm{ii})$ ${\rm E}[y(t)^{2}]<\infty$ for all $t$. $(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is nonsingular.
From Assumption 4 and (6),
$\hat{\bm{\theta}}=\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\Bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ $\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$
$=\bm{\theta}^{*}$
is obtained, and therefore $\hat{\bm{\theta}}$ is a consistent estimator of $\bm{\theta}^{*}$. In practice, the propensity score is often unknown. In such cases, the propensity score is modeled as $e^{[1]}(\bm{x};\bm{\alpha})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha})$, the parameter $\bm{\alpha}$ is estimated as $\hat{\bm{\alpha}}$ by the maximum likelihood estimation (MLE), and the propensity score in (7) is replaced by $e^{[1]}(\bm{x};\hat{\bm{\alpha}})$ to obtain a semiparametric DID estimator. If the propensity score model is correctly specified, i.e., it can be written as
${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$ (8)
and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ with some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
3 Covariate balancing for difference-in-difference approach
3.1 Estimation method and statistical properties
Propensity scores are often estimated by the MLE, but if the model for the propensity score is misspecified, the estimator of the ATT has bias. Therefore, we deal with robust methods for the model misspecification of the propensity score. Since doubly robust estimation incorporating estimators of outcome regressions has been addressed in Sant’Anna and Zhao (2020) and Ning et al. (2020a), here we construct doubly robust estimators by incorporating covariate balancing (Imai and Ratkovic 2014). Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and consider the following moment conditions:
$\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{x}\bm{x}^{{\rm T}}\bigg{]}=\bm{{\rm O}},$
$\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{x}\bm{x}^{{\rm T}}\bigg{]}=\bm{{\rm O}}.$
From these equations, ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$ is obtained, which means that the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting by the propensity score. In the original covariate balancing (Imai and Ratkovic 2014), the first-order moments of the covariates are balanced, that is, the $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$. In this article, the second-order moment of the covariates is balanced.
The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix and defining
$\bm{h}(\bm{d},\bm{x};\bm{\alpha})\equiv({\rm vech}\{\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}},{\rm vech}\{\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}})^{{\rm T}},$
the moment condition can be written as
${\rm E}\{\bm{h}(\bm{d},\bm{x};\bm{\alpha})\}=\bm{0}.$ (9)
[…]
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0},$ (10)
[…]
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$ (11)
is used. Specifically, $\hat{\bm{\alpha}}$ can be obtained by the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
Here, we use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\bm{W}_{n}\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by the Covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv 1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$. Defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}$, $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
$\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}_{n}(\hat{\bm{\alpha}}^{\rm CB})=\bm{0}.$ (12)
Here, $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix, the simplest of which is the identity matrix, but the optimal matrix, $\{1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimation of $\bm{\alpha}$. However, this optimal matrix is not always practical because it may be too close to the degenerate matrix depending on the data, causing instability in estimation.
Remark 1.
The idea of using constraints that approximately satisfy the empirical version of the moment condition to estimate the parameters of the propensity score has also been used in Ning et al. (2020b). In that article, dealing with the case of high dimensional covariates, they propose a doubly robust estimation method that combines the semiparametric estimation method of ATT with the sparse estimation method. It is also assumed that the outcome is expressed in a linear model of covariates, and to improve the convergence speed of the Horvitz-Thompson estimator, the weak covariate balancing property:
$\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\biggr{)}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0$
is included as a constraint in estimating the parameter of the propensity score.
Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994), which are necessary to describe the properties of the proposed estimation method.
Assumption 5 (Consistency).
$(\mathrm{i})$ The $\bm{W}_{n}$ converges to $\bm{W}$, which is a semi-positive definite matrix, in probability. Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$. $(\mathrm{ii})$ The $\bm{\alpha}$ is the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$. $(\mathrm{iii})$ The $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable with probability 1 for each $\bm{\alpha}\in\bm{A}$, that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is differentiable with probability 1. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\bm{A}}\|\bm{h}(\bm{d},\bm{x};\bm{\alpha})\|]<\infty$.
Assumption 6 (Asymptotic normality).
$(\mathrm{i})$ $\bm{\alpha}^{\dagger}$ is interior of $\bm{A}$. $(\mathrm{ii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$, with probability approaching one. $(\mathrm{iii})$ ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]<\infty$. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$. $(\mathrm{v})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is nonsingular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
Under the above regularity conditions, the proposed method has the following double robustness.
Theorem 1.
Replacing the propensity score in (7) by $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$, we name the covariate balancing for DID (CBD) estimator and denote it by $\hat{\bm{\theta}}^{\rm CBD}$. If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in the outcome follows a linear model of covariates, that is, if
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{x}^{{\rm T}}\bm{\beta}^{[k]*}$ (13)
holds, then under Assumptions 1-6, this $\hat{\bm{\theta}}^{\rm CBD}$ will converge to $\bm{\theta}^{*}$ in probability. Namely, proposed estimator has double robustness.
Proof.
From Assumption 5, $\hat{\bm{\alpha}}^{\rm CB}$ is consistent with $\bm{\alpha}^{\dagger}$. By expanding $\bm{h}_{n}(\hat{\bm{\alpha}})$ around $\bm{\alpha}^{\dagger}$ and applying Assumption 6 and (12), it is shown to be
$\sqrt{n}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})$
=$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (14)
$\stackrel{{\scriptstyle\rm d}}{{\to}}{\rm N}(\bm{0},\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{\Omega}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}^{-1}),$
where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}^{{\rm T}}(\bm{d},\bm{x};\bm{\alpha})]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$,
$\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$
=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1)$ (15)
is also obtained. If the propensity score model is correct and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator for $\bm{\theta}^{*}$. On the other hand, even if the model for the propensity score is not correct, when the change in outcome follows a linear model of covariates, that is, when (7) holds,
$a(\bm{x})={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]=\bm{x}^{{\rm T}}(\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*})$
is obtained from (3). Then, from (6),
$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}$
holds. Therefore,
${\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg{\{}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]-\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[0]}\mid\bm{x}_{i},d^{[0]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{\}}\biggr{]}$
$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\biggl{\{}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\biggl{(}\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\biggr{\}}\biggr{]}$
$={\rm E}\biggl{[}\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}\biggr{]}$
$=\bm{0}$
is obtained. Here, the last equality holds from the moment condition (9). From the above, when the model of the propensity score is correctly specified, or when the change in outcomes follows a linear model of covariates,
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}={\rm O}_{\rm P}{(\sqrt{n})}$
holds in (15) and $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
Remark 2.
Here, we discuss the relationship between the proposed method and recent methods. Sant’Anna and Zhao (2020) proposed a doubly robust estimating method in the study of estimating ATT by SDID, in which the propensity score and outcome are estimated by parametric models, respectively. If one of the models is correctly specified, their proposed estimator is consistent with the ATT. On the other hand, our approach does not estimate the parameters of the model for the outcomes, although the estimator is doubly robust because of the use of covariate balancing. In addition, the second-order moment of the covariates is balanced across groups targeting the treatment group population.
3.2 Numerical experiments
The robustness of the proposed estimating method against model misspecification is examined by simulation. The settings for data generation follow
$x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+\alpha^{*}x_{2})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\beta^{*}x_{1}+\epsilon^{[1]}.$
Throughout this article, all random numbers are generated independently. The parameter $\beta^{*}$ represents the magnitude of the contribution of the covariate $x_{1}$ to the causal effect. The propensity score is estimated using only $x_{1}$, that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution of $x_{2}$ to the true propensity score, which is not included in the model for estimating the propensity score. Note that when $\alpha^{*}\neq 0$, we are considering a model misspecification in which $x_{2}$ should be included in the model. Be that as it may, since (7) holds, the proposed method is theoretically guaranteed to be consistent. Under this setting, we compare the proposed method with a naive method estimating the propensity score by MLE, and confirm whether the proposed method is robust to the misspecification of the propensity score model by numerical experiments. The number of repetitions is 3000 in all numerical experiments of this article.
In Table 1, the true values of the ATT, the mean of the estimates and their empirical confidence intervals are given for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$. With increasing $\alpha^{*}$, which represents the degree of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is no longer included in the confidence interval. Regarding the proposed method, it is found that the approximation accuracy is considerably high when the weighting matrix in the generalized method of moments is the identity matrix. On the other hand, when the weighting matrix is an optimal matrix, the accuracy is often comparable to that of the identity matrix, but the accuracy is similar to that of the naive method in some cases. Since there are settings in the simulation where the optimal matrix is close to degenerate, this may be attributed to the fact that the estimation of the $\theta$ is not stable. Regardless of whether or not (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate ATT without bias.
Table 1: Comparison of the estimator of the ATT when the propensity score is given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix. Est is the average of the estimators obtained for each simulation, and Conf Int is its empirical 95% confidence interval.
CBD-id CBD-opt MLE
$\beta^{*}$ $\alpha^{*}$ $n$ True Est Conf Int Est Conf Int Est Conf Int
200 0.09 0.09 [--0.12, 0.30] 0.17 [--0.20, 0.51] --0.00 [--0.22, 0.21]
1.0 400 0.08 0.09 [--0.07, 0.24] 0.17 [--0.14, 0.45] --0.01 [--0.17, 0.14]
600 0.08 0.08 [--0.03, 0.21] 0.19 [--0.13, 0.12] --0.01 [--0.13, 0.12]
200 0.10 0.09 [--0.13, 0.32] 0.13 [--0.10, 0.37] --0.14 [--0.43, 0.15]
0.1 2.0 400 0.10 0.10 [--0.08, 0.25] 0.13 [--0.04, 0.29] --0.13 [--0.34, 0.06]
600 0.09 0.10 [--0.04, 0.24] 0.13 [--0.01, 0.27] --0.12 [--0.30, 0.03]
200 0.09 0.11 [--0.15, 0.39] 0.12 [--0.14, 0.37] --0.31 [--0.80, 0.10]
3.0 400 0.09 0.11 [--0.07, 0.29] 0.13 [--0.06, 0.31] --0.30 [--0.60, --0.03]
600 0.10 0.11 [--0.03, 0.26] 0.12 [--0.02, 0.27] --0.30 [--0.55, --0.08]
200 0.42 0.43 [0.21, 0.65] 0.52 [0.09, 0.93] 0.31 [0.06, 0.57]
1.0 400 0.43 0.43 [0.27, 0.59] 0.53 [0.17, 0.85] 0.31 [0.15, 0.49]
600 0.43 0.43 [0.31, 0.56] 0.55 [0.20, 0.80] 0.31 [0.18, 0.45]
200 0.47 0.46 [0.23, 0.70] 0.50 [0.27, 0.75] 0.21 [--0.10, 0.51]
0.5 2.0 400 0.46 0.47 [0.29, 0.63] 0.50 [0.32, 0.67] 0.21 [--0.00, 0.41]
600 0.46 0.47 [0.33, 0.61] 0.51 [0.37, 0.64] 0.22 [0.04, 0.39]
200 0.47 0.50 [0.24, 0.78] 0.51 [0.25, 0.77] 0.05 [--0.45, 0.47]
3.0 400 0.47 0.50 [0.32, 0.69] 0.52 [0.33, 0.70] 0.06 [0.32, 0.69]
600 0.48 0.50 [0.35, 0.66] 0.51 [0.37, 0.66] 0.05 [--0.20, 0.28]
200 0.85 0.85 [0.62, 1.10] 0.97 [0.47, 1.44] 0.71 [0.44, 1.01]
1.0 400 0.85 0.85 [0.68, 1.03] 0.98 [0.53, 1.35] 0.70 [0.52, 0.91]
600 0.85 0.85 [0.72, 0.99] 1.00 [0.57, 1.31] 0.71 [0.55, 0.87]
200 0.92 0.93 [0.68, 1.19] 0.97 [0.73, 1.24] 0.63 [0.30, 0.96]
1.0 2.0 400 0.92 0.93 [0.74, 1.10] 0.97 [0.78, 1.15] 0.64 [0.40, 0.85]
600 0.92 0.93 [0.78, 1.08] 0.98 [0.83, 1.12] 0.65 [0.46, 0.83]
200 0.95 0.99 [0.71, 1.28] 1.00 [0.73, 1.27] 0.50 [--0.02, 0.93]
3.0 400 0.95 0.99 [0.80, 1.18] 1.00 [0.81, 1.20] 0.50 [0.18, 0.80]
600 0.95 0.99 [0.83, 1.15] 1.00 [0.85, 1.16] 0.50 [0.24, 0.74]
200 2.56 2.55 [2.13, 2.96] 2.77 [1.84, 3.49] 2.30 [1.82, 2.82]
1.0 400 2.56 2.55 [2.27, 2.86] 2.78 [1.96, 2.67] 2.30 [1.95, 2.67]
600 2.56 2.55 [2.33, 2.76] 2.82 [2.03, 3.32] 2.30 [2.02, 2.58]
200 2.75 2.78 [2.42, 3.18] 2.86 [2.51, 3.24] 2.35 [1.91, 2.86]
3.0 2.0 400 2.75 2.78 [2.51, 3.04] 2.85 [2.60, 3.11] 2.35 [2.01, 2.68]
600 2.75 2.78 [2.57, 3.00] 2.85 [2.66, 3.06] 2.35 [2.07, 2.63]
200 2.85 2.93 [2.55, 3.33] 2.95 [2.59, 3.32] 2.28 [1.70, 2.85]
3.0 400 2.85 2.93 [2.67, 3.21] 2.95 [2.70, 3.21] 2.28 [1.89, 2.67]
600 2.85 2.93 [2.71, 3.15] 2.95 [2.75, 3.16] 2.28 [1.97, 2.59]
4 Model selection criteria for the semiparametric difference-in-difference method when propensity score is known
4.1 Risk function and its asymptotic evaluation in the semiparametric difference-in-difference method
As mentioned in the Introduction, there are no reasonable model selection criteria even for the basic setting as treated in Abadie (2005), we first derive model selection criteria when the propensity score is known, and then develop model selection criteria when using the proposed CBD method. Specifically, using the ideas of Baba et al. (2017) and Baba and Ninomiya (2021), we derive model selection criteria as asymptotically unbiased estimators of the natural risk function corresponding to the loss function used to obtain SDID estimators. By using these criteria, the dimension of $\bm{\theta}$ is determined and conponent of $\bm{x}$ is selected. As its natural risk function, based on (4), we define a weighted risk function,
$\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
$=\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
$\ \phantom{=}-\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]\}^{2}]$
$\ \phantom{=}+2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]\}$
[…]
According to the derivation of the conventional $C_{p}$ criterion, we remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term. For simplicity, we refer to the third term as the bias when the risk is simply estimated. If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), then the third term can be written as
$2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})\}\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (17)
The second equality follows from ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$ holds from (3). Here, we use the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that used to construct $\hat{\bm{\theta}}$. Then, letting $c^{\rm limit}$ be the weak limit of the third term with expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (16). Note that we consider a risk function based on the loss function in deriving the SDID estimator, but it is also possible to consider a simple squared error.
4.2 Derivation of Model Selection Criterion
In the case of known propensity scores, we denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})=e^{[1]}(\bm{x}_{i})$, $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})=\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity. The main term of the error of the SDID estimator is given by
$\hat{\bm{\theta}}-\bm{\theta}^{*}$ $={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\varepsilon_{i}+a(\bm{x}_{i})-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}.$
Thus, (17) evaluates to
${\rm E}[c^{\rm limit}]=2\sum_{i=1}^{n}{\rm E}\bigg{[}\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j})\bm{x}_{j}\{\varepsilon_{j}+a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}\}\bigg{]}.$
This expectation for the case in which $i\neq j$ is the product of the expectations for $i$ and $j$ from the independence among samples. The expected value for $i$ can be calculated as
${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]={\rm E}_{\bm{x}_{i}}[{\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]=\bm{0}.$ (18)
Therefore, we have only to consider the case $i=j$. Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from 18, and the following theorem is obtained.
Theorem 2.
When the propensity score is known,
${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\}$
is obtained, where
$\sigma^{2}(\bm{x})\equiv{\rm E}[\varepsilon^{2}\mid\bm{x}]$ $={\rm E}[\{\rho(\bm{t},\bm{x})\Delta\}^{2}\mid\bm{x}]-{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}.$ (19)
While empirical evaluation of ${\rm E}[c^{\rm limit}]$ may seem easily possible, addressing the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$ must be considered. The reason is that this cannot be empirically evaluated as $1/n\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$. In this case, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ for ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$. As a result, when the propensity score is known, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}$
$+2{\rm tr}\biggl{(}\biggl{\{}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{\}}^{-1}\biggl{[}\sum_{i=1}^{n}\Bigl{\{}\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\Bigr{\}}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{]}\biggr{)}$ (20)
as the model selection criterion for the SDID method.
4.3 Extension of the QICw
In order to provide a comparator for the proposed model selection criterion, we extend the QICw (Platt et al. 2013) for the setting of this article, which is the model selection criterion for propensity score weighted estimation on missing data. First, for the term of goodness of fit, we construct it using the same ideas as for the QICw. Specifically, the log-likelihood function with no missing data is differentiated, a weighted potential outcome variable is substituted for the one whose expected value is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted for the parameter. The number of parameters is used as the penalty term. In this setting, the log-likelihood function with no missing data should be
$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2},$
and if we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, the outcome variable and the conditional expectation will be the same from (3). Then, the sum of the goodness of fit and penalty terms multiplied by the constant $2\sigma^{2}$, that is,
$\sum_{i=1}^{n}\{\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}+2\sigma^{2}p,$
we call QICw for SDID. In fact, if $\sigma^{2}(\bm{x})$ is constant in our proposed criterion, then it is equivalent to the QICw. The problem is to estimate the variance $\sigma^{2}$ of “$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$”. This estimated value cannot be obtained from the data using only the present assumptions. Since $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ holds from $y^{[0]}(0)=y^{[1]}(0)$, and from Assumption 1, the expected value of “$\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$” is the same as the expected value of “$\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$”. Since the variance must be evaluated, we further assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, respectively, and calculate that
$\sigma^{2}$ $={\rm Var}[\Delta^{[1]}\mid d^{[1]}=1]+{\rm Var}[\Delta^{[0]}\mid d^{[0]}=1]$
$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}$
From the above, we define $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, and
$\hat{\sigma}^{2}=\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}^{2}-\bigg{(}\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}\bigg{)}^{2}+\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}^{2}-\bigg{(}\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}\bigg{)}^{2}$
is used in the present study.
Remark 3.
Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (19), if all expected values do not actually depend on $\bm{x}$, then
$\sigma^{2}(\bm{x})=$ ${\rm E}\bigg{[}\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg{|}\ d^{[1]}=1\bigg{]}-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg{[}\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg{|}\ d^{[0]}=1\bigg{]}-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
$+2{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]{\rm E}[\Delta^{[0]}\mid d^{[0]}=1].$
Since $\Delta^{[1]}$ and $\Delta^{[0]}$ include fluctuations over time, it is natural that $\Delta^{[1]}\Delta^{[0]}$ should be non-negative. In other words, basically $\sigma^{2}(\bm{x})>\sigma^{2}$, and the penalty term of the proposed criterion is larger than that of the QICw.
4.4 Numerical experiments
By simulations, we evaluate the performance of the proposed criterion (20) when the propensity score is known. As data generation settings, we consider the Case 1-1,
$x_{1}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]},$
and the Case 1-2,
$x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]}.$
In both cases, $\beta^{*}$ represents the degree of contribution of the covariate $x_{1}$ to the causal effect.
In Table 2, for some $\beta^{*}$ and sample size $n$, we examine whether the second term of (20), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate. In all cases, our penalty terms are close to the true values of the bias, while the penalty terms in the QICw significantly underestimate the bias.
Table 2: Bias evaluation when the propensity score is known: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method. The penalty term of the QICw is also given as a reference.
Case 1-1 Case 1-2
$\beta^{*}$ $n$ True Proposal QICw True Proposal QICw
200  37.54  37.29 2.23  39.20  39.43  5.96
0.1 400  38.56  37.71 2.24  39.75  39.74  5.98
600  37.89  37.89 2.27  39.79  39.68  6.00
200  57.34  56.28 2.31  59.77  58.69  6.39
0.5 400  59.36  56.28 2.33  59.36  59.52  6.48
600  58.56  57.35 2.34  60.20  59.37  6.48
200  92.48  91.33 2.56 103.24  98.85  7.91
1.0 400  98.17  91.43 2.58 108.09  99.46  7.99
600  87.03  92.62 2.59  98.33  99.57  7.99
200 359.64 358.56 5.17 439.35 429.91 23.65
3.0 400 345.73 348.96 5.20 441.47 430.88 24.05
600 367.73 365.75 5.22 435.32 430.09 23.87
Next, we examine the performance of model selection using the proposed criteria. Data is generated from
$x_{1},x_{2},\ldots,x_{l}\sim{\rm Uniform}(0,2),\quad\bm{x}=(1,x_{1},x_{2},\ldots,x_{l})^{{\rm T}},$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{\beta}^{[1]*}{}^{{\rm T}}\bm{x}+\epsilon^{[1]},$
and we consider the Case 2-1,
$l=4,\quad\bm{\beta}^{[1]*}=(1,\beta^{*},\bm{0}_{3}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
the Case 2-2,
$l=4,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{2}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
and the Case 2-3,
$l=6,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{4}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})).$
Table 3 compares the proposed criterion and the QICw by selecting models using the forward selection method. As the main index to measure the goodness of the criteria, we use the empirical estimator of the risk function, specifically, the average of 3000 calculations of
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha})(\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}.$
In the Case 2-2 setting, where there are relatively few unnecessary covariates among the candidates, that is, the negative effect of false positives is less likely to appear, the QICw slightly exceeds the proposed criterion in some cases for $\beta^{*}=0.5$ and $\beta^{*}=1.0$. Otherwise, the proposed criterion outperforms the QICw, especially in settings with many covariates unrelated to the true structure such as Case 2-1 and Case 2-3, a clear difference is observed. Since the QICw underestimates the penalty term, the false positive value must be large and too many extraneous variables must be selected in all settings.
Table 3: Comparison of the proposed criterion and the QICw through model selection when the propensity score is known: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
Case 2-1 Case 2-2 Case 2-3
$\beta^{*}$ $n$ Risk TP FP Risk TP FP Risk TP FP
200 Proposal 11.05 0.50 1.29  14.07 0.61 0.62  19.01 0.63 1.15
QICw 11.29 0.53 1.47  15.94 0.90 0.95  21.74 0.90 1.73
0.1 400 Proposal 11.04 0.47 1.30  13.99 0.68 0.52  19.72 0.69 1.08
QICw 11.27 0.51 1.49  15.23 0.98 0.86  22.45 0.98 1.77
600 Proposal 10.86 0.48 1.25  14.40 0.75 0.51  19.51 0.72 1.14
QICw 11.00 0.51 1.44  15.88 1.07 0.83  22.62 1.02 1.81
200 Proposal 14.68 0.70 1.35  23.20 1.23 0.61  30.01 1.20 1.21
QICw 14.94 0.74 1.61  23.33 1.50 1.00  31.61 1.48 1.96
0.5 400 Proposal 14.42 0.83 1.29  22.95 1.60 0.60  31.82 1.55 1.14
QICw 14.71 0.86 1.58  22.48 1.78 1.01  32.67 1.75 1.93
600 Proposal 14.67 0.89 1.31  23.24 1.74 0.57  30.25 1.75 1.20
QICw 14.93 0.91 1.61  23.00 1.88 0.98  31.96 1.88 1.98
200 Proposal 21.69 0.89 1.39  37.19 1.69 0.61  48.24 1.69 1.26
QICw 22.00 0.91 1.67  36.59 1.85 1.04  49.53 1.84 2.18
1.0 400 Proposal 21.09 0.97 1.40  34.65 1.95 0.63  46.07 1.94 1.20
QICw 21.61 0.98 1.76  35.71 1.99 1.08  49.45 1.98 2.12
600 Proposal 20.74 1.00 1.35  32.16 1.99 0.60  43.66 2.00 1.24
QICw 21.30 1.00 1.72  34.38 1.99 1.07  48.02 2.00 2.19
200 Proposal 67.79 0.99 1.54 136.61 1.96 0.67 180.41 1.96 1.33
QICw 69.22 1.00 1.93 141.09 1.99 1.19 194.02 1.98 2.34
3.0 400 Proposal 66.20 1.00 1.57 133.32 2.00 1.20 175.10 2.00 1.25
QICw 67.83 1.00 1.97 141.74 2.00 1.20 191.79 2.00 2.38
600 Proposal 65.76 1.00 1.48 126.36 2.00 0.62 173.78 2.00 1.29
QICw 67.51 1.00 1.91 135.24 2.00 1.17 191.42 2.00 2.36
5 Model Selection Criteria for the CBD Method
5.1 Derivation of Model Selection Criterion
As in the case with known propensity scores in Section 4.2, we derive the model selection criteria when using the CBD method. We evaluate
$\sum_{i=1}^{n}2{\rm E}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-{\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]\}\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*})],$
which is the third term of the risk function (17). First, substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ for ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$, and then substituting the right-hand side of (15) for $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$, this third term can be written as
$2{\rm E}\bigg{[}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}$
[…]
Since $(1/n)\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges to
$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}\{e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}\}$
in probability, if we let $\bm{u}$ be the normal random vector that is the weak limit of
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i},$ (21)
then the weak limit of (17) removing the expectation is expressed as
$c^{\rm limit}=2\bm{u}^{{\rm T}}\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}\bm{u}^{{\rm T}}\}.$
In the setting where the estimator $\hat{\bm{\theta}}^{\rm CBD}$ is consistent with $\bm{\theta}^{*}$, ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ holds because the expected value of $\bm{u}$ becomes $\bm{0}$, and the asymptotic evaluation of (17), which is to be used in the information criterion, is expressed as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}{\rm Var}[\bm{u}]\}.$
Here, we write down ${\rm Var}[\bm{u}]$. As the error of the parameter $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity score can be shown to be (14), using this, (21) can be expanded as
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bm{x}_{i}\bigg{\{}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
$\ \ \ \ \ +\dfrac{\partial}{\partial\bm{\alpha}^{{\rm T}}}\bigg{(}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bigg{[}\frac{d^{[1]}_{i}-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{]}\bigg{)}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})\bigg{\}}\{1+{\rm o}_{\rm P}(1)\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
$\ \ \ \ \ \ \ \ \ \ -\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Big{]}\{1+{\rm o}_{\rm P}(1)\},$
where
$\bm{M}(\bm{\alpha},\bm{\theta})={\rm E}\biggl{[}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x};\bm{\alpha})\biggl{[}\frac{d^{[1]}-1}{\{1-e^{[1]}(\bm{x};\bm{\alpha})\}^{2}}\Delta-\bm{x}^{{\rm T}}\bm{\theta}\biggr{]}\bm{x}\biggr{]}.$
From the above, by defining
$\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\equiv{\rm Var}$ $[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}$
$-\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})],$ (22)
it can be shown to be ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
Theorem 3.
When the propensity score is estimated by the CBD method, the third term of (16) is evaluated as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}.$
As a result, when the propensity score is estimated by the CBD method, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})(\rho(\bm{d}_{i},\bm{x}_{i})(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD})^{2}+2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}$ (23)
as the model selection criterion for the SDID method, which is an asymptotically unbiased estimator of (16).
Also in this case, the penalty term depends on the true values of the parameters, thus empirical estimates are used in practice. Specifically, $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ is used instead of $\bm{L}(\bm{\alpha}^{\dagger})$ and
$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\equiv\frac{1}{n}$ $\sum_{i=1}^{n}\Bigl{[}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
$-\bm{M}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Bigr{]}^{2}$
is used instead of $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$, where
$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\biggl{(}\frac{d^{[1]}_{i}-1}{(1-e^{[1]}(\bm{x}_{i};\bm{\alpha}))^{2}}\Delta_{i}-\bm{x_{i}}^{{\rm T}}\bm{\theta}\biggr{)}\bm{x_{i}}.$
Here, the propensity score is estimated using the covariate balancing approach, but the information criterion can be derived in the same way when the MLE is used (see Supplementary material).
5.2 Numerical experiments
Through simulations, we examine the accuracy of the approximation of the penalty term and the performance of the model selection criterion (23) based on Theorem 3 for the proposed CBD method. As the weighting matrix in the generalized method of moments, we use the identity matrix based on the results of Section 4.4. But even when an optimal matrix is used as the weighting matrix, the accuracy of the approximation of the penalty term and the performance of the model selection are comparable to those obtained when the identity matrix is used (see Supplementary Material).
In Table 4, data are generated as in Case 1-1 and Case 1-2 of Section 4.4 to verify whether the second term of (23), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate for some parameter values $\beta^{*}$ and sample size $n$. In all cases, the penalty term of the proposed criterion provides values much closer to the bias than the penalty term of the QICw, and its approximation accuracy is confirmed to be high. Conversely, in all cases, the penalty term of the QICw significantly underestimates the bias.
Table 4: Bias evaluation when the propensity score is estimated by the CBD method with identity matrix as the weighting matrix: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method. The penalty term of the QICw is also given as a reference.
Case 1-1 Case 1-2
$\beta^{*}$ $n$ True Proposal QICw True Proposal QICw
200  7.53  7.35 2.23 17.61 16.98  5.94
0.1 400  7.19  7.31 2.24 17.88 17.55  5.99
600  7.09  7.34 2.27 17.10 17.35  5.98
200  8.21  8.04 2.31 19.42 18.76  6.38
0.5 400  7.79  7.89 2.33 19.51 19.62  6.48
600  7.75  7.88 2.34 18.94 19.25  6.47
200  9.47  9.56 2.58 23.20 22.67  7.92
1.0 400  9.04  9.25 2.58 23.07 23.18  8.00
600  8.96  9.14 2.59 22.62 22.98  8.00
200 19.10 19.22 5.17 54.55 54.82 23.62
3.0 400 18.36 18.58 5.20 53.71 54.81 24.08
600 17.88 18.49 5.22 53.55 54.64 23.90
In Table 5, data are generated as in Case 2-1, Case 2-2 and Case 2-3 of Section 4.4 and the proposed criterion and the QICw are compared by model selection using the forward selection method. The main index of the goodness of the model selection criteria is again the empirical estimate of the risk. Unlike the results in Section 4.4, where the propensity score is known, the proposed criterion always outperforms the QICw, even in Case 2-2, where relatively few covariates are unnecessary among the candidates. On the other hand, as in Case 2-1 and Case 2-3, the proposed criterion is particularly superior when the candidates include many covariates that do not contribute to the true structure, which is similar to Section 4.4. The penalty term is underestimated in the QICw, therefore, too many extraneous variables must be selected.
Table 5: Comparison of the proposed criterion and the QICw through model selection when the propensity score is estimated using the CBD method with identity matrix as the weighting matrix: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
Case 2-1 Case 2-2 Case 2-3
$\beta^{*}$ $n$ Risk TP FP Risk TP FP Risk TP FP
200 Proposal 7.33 0.21 0.53 12.38 0.41 0.33 16.49 0.38 0.74
QICw 10.32 0.48 1.55 16.96 0.91 0.92 22.57 0.84 1.79
0.1 400 Proposal 7.30 0.21 0.49 12.06 0.43 0.33 17.02 0.46 0.69
QICw 10.13 0.47 1.46 15.75 0.96 0.91 22.47 0.96 1.74
600 Proposal 6.88 0.20 0.46 12.01 0.46 0.34 16.62 0.50 0.66
QICw 9.57 0.50 1.43 15.11 0.96 0.91 21.91 0.98 1.79
200 Proposal 10.73 0.51 0.55 20.12 1.12 0.34 26.11 1.11 0.72
QICw 13.12 0.77 1.65 21.56 1.55 1.01 30.24 1.51 2.01
0.5 400 Proposal 10.64 0.72 0.50 20.34 1.52 0.33 26.36 1.54 0.69
QICw 12.87 0.88 1.61 20.47 1.80 1.02 29.42 1.80 1.97
600 Proposal 10.02 0.84 0.48 17.98 1.79 0.33 24.29 1.77 0.66.
QICw 12.05 0.97 1.60 19.15 1.92 1.03 28.46 1.92 2.01
200 Proposal 13.83 0.86 0.54 25.04 1.78 0.33 34.29 1.80 0.72
QICw 17.55 0.95 1.78 28.00 1.92 1.10 41.46 1.93 2.15
1.0 400 Proposal 12.44 0.98 0.50 22.03 1.97 0.33 31.09 1.97 0.67
QICw 16.82 1.00 1.76 26.42 2.00 1.10 40.19 2.00 2.16
600 Proposal 11.33 1.00 0.48 20.06 2.00 0.31 28.79 2.00 0.63
QICw 15.91 1.00 1.76 25.28 2.00 1.11 39.30 2.00 2.17
200 Proposal 34.16 1.00 0.52 60.56 2.00 0.31 95.78 2.00 0.68
QICw 48.77 1.00 1.94 81.22 2.00 1.17 136.38 2.00 2.37
3.0 400 Proposal 31.96 1.00 0.51 59.75 2.00 0.31 91.38 2.00 0.65
QICw 46.25 1.00 1.95 80.71 2.00 1.18 132.43 2.00 2.36
600 Proposal 29.56 1.00 0.49 57.25 2.00 0.31 87.98 2.00 0.61
QICw 44.02 1.00 1.95 78.45 2.00 1.18 130.44 2.00 2.34
6 Real data analysis
The LaLonde dataset is treated in LaLonde (1986) and included in the R package Matching version 4.10.14 (Sekhon 2011). The group that took the U.S. job training program in 1976 is denoted as $t=1$ and the group that did not take the program is denoted as $t=0$. The difference in annual income in 1978 after the training for each group is estimated via the ATT. Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are treated as covariates, i.e. $p=7$, and the outcome variable is the difference between income in 1978 (re78) and income in 1974 (re78). The sample size is $N=445$. The objective of this section is not to extract information from this data, but to examine whether the proposed method makes a significant difference from existing methods, and therefore, a sufficiently small number of samples are assumed. Specifically, for a natural number $k$, the $23k-$ 2 rows of data are analyzed in Block 1, $13k-$ 1 rows in Block 2, and $3k$ rows in Block 3, respectively.
Regarding the estimation, we first estimate the propensity score using the CBD method with the identity matrix as the weighting matrix. Then, assuming that the ATT can be written as a linear sum of covariates, we estimate its coefficients by the SDID method. Table 5 selects covariates using the proposed criterion or the QICw with the forward selection method and estimates the coefficients of the selected variables. The QICw selects all variables in any Block, while the proposed criterion does not select some variables. In particular, the difference between the models selected for Block 1 and Block 3 is remarkable. Because we do not know the true structure of the real data, it is impossible to judge the superiority or inferiority of the two criteria, however, we can see that the difference between them is quite large. In a setting such as this article, it is essential to use model selection criteria that are theoretically valid rather than intuitive ones.
Table 6: Comparison of the proposed criterion and the QICw when propensity scores are estimated using the CBD method with the identity matrix as the weighing matrix, through model selection experiments on LaLonde data. The numbers with exponential notation are the estimated coefficients for each variable, and a 00 means that the variable was not selected in the model selection.
Block intercept age educ re74
1 Proposal --2.14E+3 0E+0 0E+0 0E+0
QICw --1.37E+4 5.29E+2 4.43E+1 --2.01E-1
2 Proposal 1.59E+4 0E+0 1.08E+3 2.54E-1
QICw 2.38E+4 --1.91E+2 8.22E+2 2.73E-1
3 Proposal --8.50E+2 0E+0 0E+0 0E+0
QICw --8.51E+3 6.13E+0 3.35E+2 2.56E-1
Block black hisp married nodegr
1 Proposal 0E+0 0E+0 0E+0 0E+0
QICw 3.55E+2 --8.19E+3 --9.41E+2 --1.30E+3
2 Proposal --1.86E+4 --2.34E+4 4.02E+3 --1.26E+4
QICw --1.85E+4 --2.42E+4 4.88E+3 --1.34E+4
3 Proposal 0E+0 --1.03E+4 0E+0 0E+0
QICw 3.84E+3 --6.71E+3 2.92E+3 --5.90E+2
7 Discussion
In this article, we have proposed the CBD method, a SDID method incorporating covariate balancing. In order to ensure the doubly robustness, i.e., that the ATT can be consistently estimated if the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly assumed, we have shown that it is sufficient to balance the distribution up to the second-order moment of the covariates across groups with the treatment group as the target population. This fact is important to note because covariate balancing usually balances first-order moment. We also have developed model selection criteria for the CBD method as an asymptotically unbiased estimator of the risk based on the loss function used for the estimation. Specifically, we have derived model selection criteria using a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference, and we have shown that the penalty term is considerably different from twice the number of parameters that often appears in AIC-type model selection criteria. Since there have been no model selection criteria for even the basic SDID method, we have developed the model selection criteria not only for the CBD method, but also for when the propensity score is known or the propensity score is estimated by the MLE. Numerical experiments showed that the proposed methods estimate ATT robustly compared to the method that estimates propensity scores by the MLE, and that the proposed criteria clearly reduce the risk targeted by the SDID method compared to the intuitive generalization of the existing information criterion to the SDID method. In addition, the real data analysis has confirmed that there were large differences between the results of the proposed method and the existing methods.
Although this article deals with a linear model of covariates as a simple working model for the outcome, it should be noted that the results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method. In fact, Hazlett (2020) treats the function $\phi(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in the reproducing kernel Hilbert space as $q$ may be larger than $p$ or $n$. And they assume that the outcome is represented by a linear model of $\phi(\bm{x})$, the feature map of $\bm{x}$, which in our setting can be written as
${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\theta}.$ (24)
Then, by using
$\sum_{i=1}^{n}\bigg{\{}\frac{d^{[1]}_{i}\phi(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\phi(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg{\}}=0$
as the moment condition for estimating the propensity score, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear. In the setting of this article, we also consider the situation (24) and when making the estimation doubly robust, we assume that the change in the outcomes follows
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\phi(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semidefinite symmetric matrix whose $(l,m)$ component is $k(x_{l},x_{m})$. If we estimate the parameter $\bm{\alpha}$ that satisfies
$\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{K}\bigg{]}=\bm{{\rm O}},$
$\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{K}\bigg{]}=\bm{{\rm O}},$
and then estimate $\bm{\theta}$ or derive model selection criteria based on it, that would be a natural extension of the proposed method in this article.
There are other valuable directions for extension. For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) propose a DID method in a more flexible setting than the previous designs, where there are multiple time points and the treatment may be initiated at any of them. Moreover, De Chaisemartin and D’haultfœuille (2023) develops a DID method when there are more than two groups, and Richardson et al. (2023) proposes the generalized DID to estimate causal effects under alternative assumptions to the parallel trend assumption. While the outcome in the basic DID method is continuous, Tchetgen et al. (2024) deals with the case where the response follows a generalized linear model. Thus, the DID method has been theoretically developed in accordance with real problems, and extensions of the proposed method to these problems are recent challenges. In econometrics, the synthetic control method (Abadie et al. 2010) is a major method for causal inference as well as the DID method, and a method combining it with the DID method has also been developed as synthetic DID (Arkhangelsky et al. 2021). In this method, the objective of the information criterion is what to synthesize and with what weights, and the proposed method cannot necessarily be extended in a similar way, but it is an important topic. Furthermore, propensity score analysis with high-dimensional confounding variables is also a hot topic (e.g., Belloni et al. 2017, Chernozhukov et al. 2018, Ning et al. 2020b), and even in that analysis, robustly estimating the influence of covariates that directly affect the causal effect and selecting covariates that truly affect the causal effect is an essential task to ensure the validity of the analysis. The asymptotic properties of the estimator of causal effects tend to be obtained even when the covariates are high-dimensional as well, and it is surely possible to extend our proposed method for these analyses.

### v3

@@
 Abstract
-In causal inference, remarkable progress has been made in difference-in-differences (DID) approaches to estimate the average effect of treatment on the treated (ATT).
+Remarkable progress has been made in difference-in-differences (DID) approaches to causal inference that estimate the average effect of a treatment on the treated (ATT).
 Of these, the semiparametric DID (SDID) approach incorporates a propensity score analysis into the DID setup.
 Supposing that the ATT is a function of covariates, we estimate it by weighting the inverse of the propensity score.
-As one method to make the estimation robust to the propensity score modeling, we incorporate covariate balancing.
-Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator has doubly robustness.
-In addition to the estimation, model selection is also addressed.
+In this study, as one way to make the estimation robust to the propensity score modeling, we incorporate covariate balancing.
+Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator is doubly robust.
+In addition to the estimation, we also address model selection.
 In practice, covariate selection is an essential task in statistical analysis, but even in the basic setting of the SDID approach, there are no reasonable information criteria.
-Therefore, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation.
-As a result, we show that a penalty term is derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria.
-Numerical experiments show that the proposed method estimates the ATT robustly compared to the method using propensity scores given by the maximum likelihood estimation (MLE), and that the proposed criterion clearly reduces the risk targeted in the SDID approach compared to the intuitive generalization of the existing information criterion.
-In addition, real data analysis confirms that there is a large difference between the results of the proposed method and the existing method.
-Keywords: Covariate balancing, Difference-in-differences, Information criterion, Model selection, Propensity score, Semiparametric approach, Statistical asymptotic theory
+Here, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation.
+We show that a penalty term can be derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria.
+Keywords: Average effect of treatment on the treated, Doubly robustness, Information criterion, Propensity score analysis, Statistical asymptotic theory
 1 Introduction
@@
 In the basic setting, there are two groups: a treatment group receiving a specific treatment and a control group receiving no treatment or other treatment.
-The pre and post-treatment outcomes of each group are measured and the differences in the outcomes are compared between the groups to estimate the average treatment effect on the treated (ATT).
-Under the parallel trend assumption, simple estimation is valid.
-More recently, Abadie (2005) dealt with inference under more naturalized assumptions using information from the covariates, specifically the parallel trend assumption conditioned on the covariates.
-Then, as a method that does not necessarily require modeling of the covariates to the outcome, they propose the semiparametric DID (SDID) method, which estimates ATT by weighting the inverse of the propensity score.
-With or without conditionality, the parallel trend assumption is a key assumption in the DID.
-This is the assumption that the mean outcomes of a treatment group and a control group would have been similar over time if there had been no treatment.
-On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g., Robins et al.
-1994 and Li et al.
-2018).
+The pre- and post-treatment outcomes of each group are measured and the change in the outcome for the control group is subtracted from the change in the outcome for the treatment group to estimate the average treatment effect on the treated (ATT).
+Under the parallel trend assumption, an estimation in the DID approach is valid.
+Recently, Abadie (2005) dealt with inference under a more natural assumption that involves using information from the covariates, specifically a parallel trend assumption conditioned on the covariates.
+They further proposed a semiparametric DID (SDID) approach that does not necessarily require modeling of the covariates to the outcome and that estimates the ATT by weighting the inverse of the propensity score.
+With or without conditionality, the parallel trend assumption is key to the DID approach.
+This assumption is one where the mean changes over time in the potential outcomes of the treatment group and the control group would have been similar if there had been no treatment.
+On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g.
+Robins et al.
+1994).
 This is the assumption that the potential outcome variable and the treatment assignment variable, which represents the received treatment, become independent conditioned on the covariates.
-Although neither assumption is verifiable, the validity of the parallel trend assumption can be examined by visual inspection when we have observations over many time points.
-And the parallel trend assumption naturally holds if it can be assumed that there is no change over time in the expected value of the outcome without the treatment.
-Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such desirable conditions is the advantage that the DID gained in exchange for collecting pre- and post-treatment outcomes.
-The methodological development of the DID has been remarkable in recent years because of such advantages and the increasing amount of available observational study data (see Roth et al.
-2023 for a comprehensive review of the DID in recent years).
-However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified, but if the model is incorrect, the resulting estimator has a bias.
-As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome, and proposes the estimator for the ATT that combines the results of each estimator.
-Their proposed estimator is consistent if either a propensity score or outcome regression working model is correctly specified, that is, the estimators have doubly robustness.
-In this article, as with Abadie (2005), we assume that ATT is a function of covariates and address the semiparametric estimation problem for that function.
-We assume two regression models as above and give double robustness to the estimation, but unlike the Sant’Anna and Zhao (2020), we incorporate covariate balancing (Imai and Ratkovic 2014), a method that uses propensity scores such that a certain moment condition is satisfied.
-This method actually estimates the model for the treatment assignment, while assuming, but not estimating, the model for the outcome.
-Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either one of the regression models is correctly specified.
-One point to note is that, in order to satisfy this favorable asymptotic property in DID, the moment condition is constructed in a different way from the usual one.
-Incidentally, incorporating covariate balancing leads to the estimation of causal effects via weights that always balance the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014).
-We call the proposed methodology the covariate balancing for DID (CBD) method.
-Obviously, model selection is an essential task in the DID as well, in a series of data analysis tasks.
-For example, in estimating the effect of union membership on wages, Houngbedji (2016) treats a linear model with workers’ demographics for causal effects, and explores factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value.
-In other words, the issue of covariate selection is addressed.
-However, for the information criterion that is the basic tool for model selection, there are no reasonable ones, even for the basic setting of the SDID.
-Thus, we derive the information criterion for the basic SDID and CBD methods as an asymptotically unbiased estimator of the risk based on the loss function used in the SDID estimation.
-Note that traditional methods such as GIC (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which is a random variable.
-Since the SDID method places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment condition when deriving the information criterion.
-This causes some difficulties.
-As a result, we derive penalty terms that are significantly different from twice the number of parameters, which often appear in AIC-type information criteria including GIC.
-In summary, the aims of this article are to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID method, and to derive model selection criteria for the SDID method, for which no reasonable model selection criteria exist yet.
-The organization of this article is as follows.
-First, in section 2, the model and estimation target are specified, and the SDID method proposed in Abadie (2005) is introduced.
-In section 3, we propose the CBD method and give a theorem showing its doubly robustness.
-Numerical experiments show that the estimation is robust against model misspecification.
-In section 4, when the propensity score is known, we derive the model selection criterion as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID method.
-Then, we also give a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid.
-As a comparison, we derive the model selection criterion the QICw, which naturally extends the idea of Platt et al.
-(2013) to the SDID method.
-Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of estimating the risk, and also verify its performance compared to the QICw from the viewpoint of risk minimization.
-Section 5 summarizes the results for the CBD method in the same manner as Section 4.
-In other words, model selection criteria are derived, theorems are given to support their validity, and their performance is evaluated through numerical experiments.
-Section 6 reports the results of estimating causal effects and model selection for the real data in LaLonde (1986).
-Finally, Section 7 discusses the future development of the proposed method.
+Although neither assumption is verifiable, to evaluate the validity of the parallel trend assumption, visual inspection is useful when observations over many time points are available.
+Moreover, the parallel trend assumption naturally holds if it can be supposed that there is no change over time in the expectation of outcomes without treatment.
+Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such a desirable assumption is the advantage that the DID approach gains in exchange for collecting pre- and post-treatment outcomes.
+The methodological development of the DID approach has been growing because of the above-mentioned advantages and the increasing amount of available observational study data (see Roth et al.
+2023 for a comprehensive review of recent studies on the DID approach).
+However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified.
+However, the estimator will have a bias if the model is misspecified.
+As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome variable, and proposes an estimator for the ATT derived from the two models.
+Their estimator is consistent if either the treatment assignment model or the outcome regression model is correctly specified, that is, the estimator is doubly robust.
+In this article, as with Abadie (2005), we suppose that the ATT is a function of covariates and address the semiparametric estimation problem for that function.
+From the estimated function, the heterogeneity of the ATT, that is, the conditional ATT can be evaluated.
+Two regression models as above are considered, and in order to give double robustness to the estimation, we incorporate covariate balancing (e.g.
+Egel et al.
+2008, Imai and Ratkovic 2014 and Fan et al.
+2022), a method using propensity scores such that a certain moment condition is satisfied.
+This method actually estimates the treatment assignment model, but only supposes an outcome model and does not estimate it.
+Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either of the regression models is correctly specified.
+One point to note is that, in order for the SDID approach to have this favorable asymptotic property, the moment condition has to be constructed in a different way from the usual one, which is due to considering the conditional ATT.
+Also, the incorporation of covariate balancing leads to an estimation via weights that always balances the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014).
+We call the proposed method covariate balancing for DID (CBD).
+Obviously, model selection is also an essential task in using the DID approach in a series of data-analysis tasks.
+Specifically, by selecting the function of covariates for the ATT addressed in this paper, it becomes possible to evaluate the heterogeneity of the ATT.
+There have been various important studies conducted in this area.
+For example, in a study estimating the effect of union membership on wages, Houngbedji (2016) treated the workers’ demographics with a linear model for causal effects and explored factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value.
+In other words, they addressed the issue of covariate selection.
+However, there is no satisfactory information criterion, a basic tool for model selection, even for the basic setting of SDID.
+To remedy this situation, we will derive an information criterion for a basic SDID approach and the CBD method that is an asymptotically unbiased estimator of the risk based on the loss function used in the estimation.
+Note that traditional information criteria such as the generalized information criterion (GIC) (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which are random variables.
+Since the SDID approach places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment assumption when deriving the information criterion, which causes some difficulties.
+As a result, we will derive a penalty term that is considerably different from twice the number of parameters, which often appears in AIC-type information criteria including the GIC.
+In summary, the aim of this paper is to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID approach and to derive a model selection criterion for the SDID approach, for which no reasonable model selection criteria exist yet.
+Its organization is as follows.
+Section 2 describes the model and the estimation target and introduces the SDID approach that was proposed in Abadie (2005).
+In Section 3, we propose the CBD method and give a theorem showing its double robustness.
+Numerical experiments indicate that the estimation is robust against model misspecification.
+In Section 4, we derive a model selection criterion for when the propensity score is known as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID estimation.
+Then, we also prove a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid.
+As a comparison, we derive a model selection criterion that naturally extends the idea of QICw in Platt et al.
+(2013) to the SDID approach.
+Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of risk estimation and show its superior performance to the QICw from the viewpoint of risk minimization.
+Section 5 summarizes the results for the CBD method in the same way as Section 4.
+In other words, a model selection criterion is derived from the naturally defined risk, a theorem is given to support the validity of its penalty term, and its performance is evaluated through numerical experiments.
+Section 6 reports the results of the estimation and model selection for the real data in LaLonde (1986).
+Finally, Section 7 discusses future developments of the proposed approach.
 2 Preparation
@@
 Suppose that there are two groups, a treatment group and a control group, and that outcomes are collected before and after the possible treatment time point, respectively.
-We treat the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome of the treatment group, $y^{[0]}(t)$ the outcome of the control group, $t=0$ the outcome before treatment, and $t=1$ the outcome after treatment.
-In practice, only one of the outcomes of the treatment or control group can be obtained.
-Preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if assigned to the control group, the observed outcome collected at each time point can be expressed as
+We consider the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome variable of the treatment group, $y^{[0]}(t)$ the outcome variable of the control group, $t=0$ before treatment and $t=1$ after treatment.
+In practice, only one of the outcomes of the treatment group or control group can be obtained.
+By preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if the sample is assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if the sample is assigned to the control group, the observed outcome variable collected at each time point is expressed as
 $y(t)=d^{[1]}y^{[1]}(t)+d^{[0]}y^{[0]}(t).$
-Letting the change in the outcome variable at two time points in the treatment group be $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group be $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change can be expressed as
+Denoting the change in the outcome variables at two time points in the treatment group as $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group as $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change is expressed as
 $\Delta=d^{[1]}\Delta^{[1]}+d^{[0]}\Delta^{[0]}.$
@@
 $a(\bm{x})\equiv{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1],$ (1)
-we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, we suppose the working model to be a linear model $\bm{x}^{{\rm T}}\bm{\theta}$.
+we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, set the working model to be a linear one $\bm{x}^{{\rm T}}\bm{\theta}$.
 Then, we estimate $\bm{\theta}$ and give the ATT by taking the expectation with respect to $\bm{x}$ in the treatment group.
-Now, we give the optimal value of $\bm{\theta}$ as
+Let us define the optimal value of $\bm{\theta}$ as
 $\bm{\theta}^{*}\equiv\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[\{{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}\mid d^{[1]}=1].$ (2)
-Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated by the SDID method using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and thus the ATT can be consistently estimated.
+Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated with the SDID approach by using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and therefore the ATT can be consistently estimated.
 The assumptions are as follows.
 Assumption 1 (Conditional parallel trend).
+It holds that
 ${\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[0]}=1],$
-that is, for the expected value conditioned on $\bm{x}$ of the change in outcome over time without the treatment, the values for the treatment and control groups are the same.
+which means that, under the expectation conditioned on $\bm{x}$, the change in outcomes without treatment is the same regardless of whether it is the treatment group or the control group.
 Assumption 2 (Common shocks).
-Events occurring between the time points of outcome collection before and after the treatment have the same impact on outcomes in both the treatment group and the control group.
+Events that occur between the time points of outcome collection before and after treatment have the same impact on outcomes in both the treatment group and the control group.
 Assumption 3 (Positivity).
-$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1,$
-that is, the propensity score is not 00 or $1$.
-Abadie (2005) showed that using the weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$, then
+It holds that
+$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1;$
+that is, the propensity score is not 0 or $1$.
+Abadie (2005) showed that, by using weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$,
 ${\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]$
-$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]$
-$={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]$
-$=a(\bm{x})$ (3)
+$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=a(\bm{x})$ (3)
 holds.
-Assumption 1 is used in the second equation.
-This equation means that by multiplying weights based on the propensity scores to the observed values of change, the conditional ATT can be estimated via DID.
-Then, from (2) and (3), it is shown that the optimal solution of $\bm{\theta}$ can be expressed as
+Assumption 1 is used in the second equality.
+This equation means that, by multiplying the observed values of change with weights based on propensity scores, the conditional ATT can be estimated via the DID approach.
+Then, from (2) and (3), the optimal solution of $\bm{\theta}$ can be expressed as
 $\bm{\theta}^{*}=\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[e^{[1]}(\bm{x})\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}].$ (4)
-From this equation,
-${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$ $={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
-$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
-$=\bm{0}$ (5)
-is satisfied, and therefore, the equation
+Moreover, from this equation,
+${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
+$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]=\bm{0}$ (5)
+is satisfied, and therefore,
 ${\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$ (6)
@@
 2.2 Semiparametric estimation
-Given an i.i.d. sample of size $n$, the $i$-th sample is denoted with the subscript $i$. Abadie (2005) proposed
+Here, for an independently and identically distributed set of samples of size $n$, the $i$-th sample will be denoted by the subscript $i$. Abadie (2005) proposed
 $\hat{\bm{\theta}}=\bigg\{\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ (7)
 as a natural estimator of $\bm{\theta}^{*}$ by solving (4).
-When the weights based on the propensity scores are multiplied to the differences of the observed values, a term corresponding to the difference in difference appears, thus this method is called the Semiparametric DID (SDID) estimator.
-The following regular conditions are assumed.
+When the observed values of change are multiplied by weights based on propensity scores, a term corresponding to the difference-in-differences appears; thus, this method is called a semiparametric DID (SDID) estimator.
+The following regularity conditions are assumed.
 Assumption 4.
-$(\mathrm{i})$ $\bm{\theta}^{*}$ is interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$.
+$(\mathrm{i})$ $\bm{\theta}^{*}$ is in the interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$.
 $(\mathrm{ii})$ ${\rm E}[y(t)^{2}]<\infty$ for all $t$.
-$(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is nonsingular.
-From Assumption 4 and (6),
-$\hat{\bm{\theta}}=\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\Bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ $\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
-$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]$
-$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$
-$=\bm{\theta}^{*}$
-is obtained, and therefore $\hat{\bm{\theta}}$ is a consistent estimator of $\bm{\theta}^{*}$.
-In practice, the propensity score is often unknown.
-In such cases, the propensity score is modeled as $e^{[1]}(\bm{x};\bm{\alpha})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha})$, the parameter $\bm{\alpha}$ is estimated as $\hat{\bm{\alpha}}$ by the maximum likelihood estimation (MLE), and the propensity score in (7) is replaced by $e^{[1]}(\bm{x};\hat{\bm{\alpha}})$ to obtain a semiparametric DID estimator.
-If the propensity score model is correctly specified, i.e., it can be written as
-${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$ (8)
-and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ with some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
-3 Covariate balancing for difference-in-difference approach
-3.1 Estimation method and statistical properties
-Propensity scores are often estimated by the MLE, but if the model for the propensity score is misspecified, the estimator of the ATT has bias.
-Therefore, we deal with robust methods for the model misspecification of the propensity score.
-Since doubly robust estimation incorporating estimators of outcome regressions has been addressed in Sant’Anna and Zhao (2020) and Ning et al.
-(2020a), here we construct doubly robust estimators by incorporating covariate balancing (Imai and Ratkovic 2014).
-Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and consider the following moment conditions:
+$(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is non-singular.
+From Assumption 4 and (6), we obtain
+$\hat{\bm{\theta}}=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$
+$\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
+$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}=\bm{\theta}^{*},$
+[…]
+${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$
+and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ under some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
+3 Covariate balancing for difference-in-differences approach
+3.1 Estimation method and its statistical properties
+Propensity scores are often estimated by using MLE, but if the model for the propensity scores is misspecified, the estimator of the ATT will have bias.
+Here, we propose a robust method that can deal with model misspecification.
+While a doubly robust estimation known as augmented inverse-probability-weighted estimation, which uses estimators from an outcome regression model, has been addressed in Sant’Anna and Zhao (2020) and Ning et al.
+(2020a), here we construct a doubly robust estimation by incorporating covariate balancing (e.g.
+Egel et al.
+2008, Imai and Ratkovic 2014 and Fan et al.
+2022).
+Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and let us consider the following moment conditions:
 $\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{x}\bm{x}^{{\rm T}}\bigg]=\bm{{\rm O}},$
 $\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{x}\bm{x}^{{\rm T}}\bigg]=\bm{{\rm O}}.$
-From these equations, ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$ is obtained, which means that the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting by the propensity score.
-In the original covariate balancing (Imai and Ratkovic 2014), the first-order moments of the covariates are balanced, that is, the $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$.
-In this article, the second-order moment of the covariates is balanced.
-The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix and defining
+From these conditions is obtained ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$, which means that the expectation of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expectation of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting based on the propensity scores.
+In the original covariate balancing, the first-order moments of the covariates are balanced; that is, $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$.
+In this paper, the second-order moments of the covariates are balanced.
+The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix, by letting
 $\bm{h}(\bm{d},\bm{x};\bm{\alpha})\equiv({\rm vech}\{\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}},{\rm vech}\{\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}})^{{\rm T}},$
-the moment condition can be written as
-${\rm E}\{\bm{h}(\bm{d},\bm{x};\bm{\alpha})\}=\bm{0}.$ (9)
-[…]
-$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0},$ (10)
-[…]
-$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$ (11)
-is used. Specifically, $\hat{\bm{\alpha}}$ can be obtained by the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
-Here, we use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\bm{W}_{n}\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by the Covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv 1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$.
-Defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}$, $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
+the moment conditions can be rewritten as
+${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}.$ (8)
+We will denote an $\bm{\alpha}$ satisfying (8) as $\bm{\alpha}^{\dagger}$.
+Since the propensity score model may be misspecified, $\bm{\alpha}^{\dagger}$ is not necessarily the true value $\bm{\alpha}^{*}$, but when the propensity score model is correctly specified, $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$.
+In practice, the SDID approach uses a $\hat{\bm{\alpha}}$ satisfying the empirical version of the moment conditions,
+$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0}.$ (9)
+If the moment conditions are just-identified, that is, the dimension of the moment conditions (the number of constraints) is the same as the dimension of the parameter $\bm{\alpha}$, then $\bm{\alpha}$ that satisfies (9) exists.
+On the other hand, if they are over-identified, that is, the dimension of the moment conditions is larger than the dimension of $\bm{\alpha}$, then an $\hat{\bm{\alpha}}$ that satisfies (9) does not necessarily exist.
+In that case, an $\hat{\bm{\alpha}}$ satisfying
+$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$
+is used. Specifically, $\hat{\bm{\alpha}}$ is obtained from the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
+Here, we will use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\allowbreak\bm{W}_{n}\allowbreak\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by using the covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv n^{-1}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$ and $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix.
+Upon defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}$, one sees that the estimator $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
 $\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}_{n}(\hat{\bm{\alpha}}^{\rm CB})=\bm{0}.$ (10)
-Here, $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix, the simplest of which is the identity matrix, but the optimal matrix, $\{1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimation of $\bm{\alpha}$.
-However, this optimal matrix is not always practical because it may be too close to the degenerate matrix depending on the data, causing instability in estimation.
+Its simplest candidate is the identity matrix, but the optimal matrix, $\{n^{-1}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\allowbreak\bm{h}(\bm{d}_{i},\allowbreak\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimator of $\bm{\alpha}$.
+However, this optimal matrix is not necessarily practical since it may be too close to being a degenerate matrix for some data, causing considerable instability in the estimation of $\bm{\alpha}$ and consequently $\bm{\theta}$.
 Remark 1.
-The idea of using constraints that approximately satisfy the empirical version of the moment condition to estimate the parameters of the propensity score has also been used in Ning et al.
-(2020b).
-In that article, dealing with the case of high dimensional covariates, they propose a doubly robust estimation method that combines the semiparametric estimation method of ATT with the sparse estimation method.
-It is also assumed that the outcome is expressed in a linear model of covariates, and to improve the convergence speed of the Horvitz-Thompson estimator, the weak covariate balancing property:
-$\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\biggr{)}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0$
-is included as a constraint in estimating the parameter of the propensity score.
-Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994), which are necessary to describe the properties of the proposed estimation method.
+The idea of using constraints that approximately satisfy the empirical version of the moment conditions to estimate the parameters of the propensity scores has also been examined.
+In Ning et al.
+(2020b) a doubly robust estimation method is proposed for dealing with the case of high-dimensional covariates.
+This method combines a semiparametric estimation of the ATT with sparse estimation.
+It is assumed that the outcome can be expressed by a linear model of covariates, and to speed up the convergence of the Horvitz-Thompson estimator, the weak covariate balancing property, written as
+$\frac{1}{n}\sum_{i=1}^{n}\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\bigg\}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0,$
+is included as a constraint in estimating the parameter of the propensity scores.
+Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994) that are necessary to derive the properties of the proposed estimation method.
 Assumption 5 (Consistency).
-$(\mathrm{i})$ The $\bm{W}_{n}$ converges to $\bm{W}$, which is a semi-positive definite matrix, in probability.
-Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$.
-$(\mathrm{ii})$ The $\bm{\alpha}$ is the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$.
-$(\mathrm{iii})$ The $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable with probability 1 for each $\bm{\alpha}\in\bm{A}$, that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is differentiable with probability 1.
+$(\mathrm{i})$ $\bm{W}_{n}$ converges in probability to $\bm{W}$, which is a semi-positive definite matrix.
+Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if $\bm{W}{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$.
+$(\mathrm{ii})$ $\bm{\alpha}^{\dagger}$ is in the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$.
+$(\mathrm{iii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is continuous; that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is continuous at each $\bm{\alpha}\in\bm{A}$ with probability one.
 $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\bm{A}}\|\bm{h}(\bm{d},\bm{x};\bm{\alpha})\|]<\infty$.
 Assumption 6 (Asymptotic normality).
-$(\mathrm{i})$ $\bm{\alpha}^{\dagger}$ is interior of $\bm{A}$.
-$(\mathrm{ii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$, with probability approaching one.
-$(\mathrm{iii})$ ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]<\infty$.
-$(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$.
-$(\mathrm{v})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is nonsingular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
-Under the above regularity conditions, the proposed method has the following double robustness.
+$(\mathrm{i})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is continuously differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$ with probability approaching one.
+$(\mathrm{ii})$ ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})]=\bm{0}$ and ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]\allowbreak<\infty$.
+$(\mathrm{iii})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\allowbreak\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$.
+$(\mathrm{iv})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is non-singular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
+Under these regularity conditions, the proposed method can be shown to be doubly robust, as follows.
+Here, we name the estimator obtained by replacing the propensity scores with $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$ in (7) the covariate balancing for DID (CBD) estimator and denote it as $\hat{\bm{\theta}}^{\rm CBD}$.
 Theorem 1.
-Replacing the propensity score in (7) by $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$, we name the covariate balancing for DID (CBD) estimator and denote it by $\hat{\bm{\theta}}^{\rm CBD}$.
-If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in the outcome follows a linear model of covariates, that is, if
+If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in outcomes follows a linear model of covariates, that is, if
 ${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{x}^{{\rm T}}\bm{\beta}^{[k]*}$ (11)
-holds, then under Assumptions 1-6, this $\hat{\bm{\theta}}^{\rm CBD}$ will converge to $\bm{\theta}^{*}$ in probability. Namely, proposed estimator has double robustness.
+holds, then under Assumptions 1-6, $\hat{\bm{\theta}}^{\rm CBD}$ converges in probability to $\bm{\theta}^{*}$. That is, the proposed estimator has double robustness.
 Proof.
-From Assumption 5, $\hat{\bm{\alpha}}^{\rm CB}$ is consistent with $\bm{\alpha}^{\dagger}$.
-By expanding $\bm{h}_{n}(\hat{\bm{\alpha}})$ around $\bm{\alpha}^{\dagger}$ and applying Assumption 6 and (12), it is shown to be
+[…]
 $\sqrt{n}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})$
-=$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (14)
+$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (12)
 $\stackrel{{\scriptstyle\rm d}}{{\to}}{\rm N}(\bm{0},\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{\Omega}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}^{-1}),$
-where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}^{{\rm T}}(\bm{d},\bm{x};\bm{\alpha})]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$,
+where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}(\bm{d},\bm{x};\bm{\alpha})^{{\rm T}}]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$, we also obtain
 $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$
-=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
-=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1)$ (15)
-is also obtained.
-If the propensity score model is correct and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator for $\bm{\theta}^{*}$.
-On the other hand, even if the model for the propensity score is not correct, when the change in outcome follows a linear model of covariates, that is, when (7) holds,
+$=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$ (13)
+$=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1).$ (14)
+If the propensity score model is correctly specified and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator of $\bm{\theta}^{*}$.
+On the other hand, even if the propensity score model is misspecified, when the change in outcomes follows a linear model of covariates, that is, when (7) holds, we obtain
 $a(\bm{x})={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]=\bm{x}^{{\rm T}}(\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*})$
-is obtained from (3). Then, from (6),
-$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}$
-holds. Therefore,
+from (3). Then, it follows from (6) that
+$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}.$
+Therefore, we obtain
 ${\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
-$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg{\{}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]-\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[0]}\mid\bm{x}_{i},d^{[0]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{\}}\biggr{]}$
-$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\biggl{\{}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\biggl{(}\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\biggr{\}}\biggr{]}$
-$={\rm E}\biggl{[}\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}\biggr{]}$
-$=\bm{0}$
-is obtained.
-Here, the last equality holds from the moment condition (9).
-From the above, when the model of the propensity score is correctly specified, or when the change in outcomes follows a linear model of covariates,
+$={\rm E}\bigg[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]$
+[…]
+$={\rm E}\bigg[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg[\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\bigg\}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\bigg\{\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\bigg\}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\bigg]\bigg]$
+$={\rm E}[\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}]=\bm{0}.$
+Here, the last equality holds from the moment conditions in (8).
+From the above, when the propensity score model is correctly specified, or when the change in outcomes follows a linear model of covariates, it holds that
 $\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}={\rm O}_{\rm P}{(\sqrt{n})}$
-holds in (15) and $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
+in (14), and therefore $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
 Remark 2.
-Here, we discuss the relationship between the proposed method and recent methods.
-Sant’Anna and Zhao (2020) proposed a doubly robust estimating method in the study of estimating ATT by SDID, in which the propensity score and outcome are estimated by parametric models, respectively.
-If one of the models is correctly specified, their proposed estimator is consistent with the ATT.
-On the other hand, our approach does not estimate the parameters of the model for the outcomes, although the estimator is doubly robust because of the use of covariate balancing.
-In addition, the second-order moment of the covariates is balanced across groups targeting the treatment group population.
+Let us discuss the relationship between our proposed method and the method in Sant’Anna and Zhao (2020).
+Unlike our setting, which estimates the conditional ATT, the latter directly estimates the ATT using a doubly robust approach.
+In their approach as well, the propensity scores are estimated so that the first-order moments of the distribution for the covariates are balanced.
+They also suppose a parametric model of the covariates for the change in the control group’s outcomes and estimate it, using these estimators to estimate the ATT.
+In contrast, we place importance on evaluating the heterogeneity of the ATT, including variable selection, and then set a function of the covariates for the ATT.
+We suppose a parametric model of the covariates for the change in outcomes as well, although not estimating it.
+Then, if the propensity scores are estimated so that not the first-order but the second-order moments of the distribution for the covariates are balanced, it is easily, yet somewhat unexpectedly, shown that the conditional ATT estimator is doubly robust.
 3.2 Numerical experiments
-The robustness of the proposed estimating method against model misspecification is examined by simulation. The settings for data generation follow
+Let us examine the robustness of the proposed estimation against model misspecification through simulation studies. The setting for data generation follows
 $x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+\alpha^{*}x_{2})),$
@@
 $y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\beta^{*}x_{1}+\epsilon^{[1]}.$
-Throughout this article, all random numbers are generated independently.
-The parameter $\beta^{*}$ represents the magnitude of the contribution of the covariate $x_{1}$ to the causal effect.
-The propensity score is estimated using only $x_{1}$, that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution of $x_{2}$ to the true propensity score, which is not included in the model for estimating the propensity score.
-Note that when $\alpha^{*}\neq 0$, we are considering a model misspecification in which $x_{2}$ should be included in the model.
-Be that as it may, since (7) holds, the proposed method is theoretically guaranteed to be consistent.
-Under this setting, we compare the proposed method with a naive method estimating the propensity score by MLE, and confirm whether the proposed method is robust to the misspecification of the propensity score model by numerical experiments.
-The number of repetitions is 3000 in all numerical experiments of this article.
-In Table 1, the true values of the ATT, the mean of the estimates and their empirical confidence intervals are given for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$.
-With increasing $\alpha^{*}$, which represents the degree of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is no longer included in the confidence interval.
-Regarding the proposed method, it is found that the approximation accuracy is considerably high when the weighting matrix in the generalized method of moments is the identity matrix.
-On the other hand, when the weighting matrix is an optimal matrix, the accuracy is often comparable to that of the identity matrix, but the accuracy is similar to that of the naive method in some cases.
-Since there are settings in the simulation where the optimal matrix is close to degenerate, this may be attributed to the fact that the estimation of the $\theta$ is not stable.
-Regardless of whether or not (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate ATT without bias.
-Table 1: Comparison of the estimator of the ATT when the propensity score is given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix.
-Est is the average of the estimators obtained for each simulation, and Conf Int is its empirical 95% confidence interval.
+Throughout this paper, we will suppose that all random numbers are generated independently.
+The parameter $\beta^{*}$ represents the magnitude of the contribution from the covariate $x_{1}$ to the causal effect.
+The propensity score is estimated using only $x_{1}$; that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution from $x_{2}$ to the true propensity scores, which is not included in the model for estimating the propensity scores.
+Note that when $\alpha^{*}\neq 0$, $x_{2}$ should also be included in the model, and we are doing a model misspecification because it was not included.
+Anyway, since (7) holds, the proposed estimation is theoretically guaranteed to be consistent.
+In this setting, we will compare the proposed method with a naive method estimating the propensity score by the MLE and check whether the proposed method is robust to misspecification of the propensity score model.
+Note that each of the numerical experiments described below was repeated 3,000 times.
+The true values of the ATT, the mean estimates, and their empirical confidence intervals are listed in Table 1 for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$.
+With increasing $\alpha^{*}$, which represents the magnitude of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is eventually no longer included in the confidence interval.
+Regarding the proposed method, the approximation accuracy is considerably high when the identity matrix is used as the weighting matrix in the generalized method of moments.
+On the other hand, when the optimal matrix is used as the weighting matrix, although the accuracy is often comparable to that of using the identity matrix, in some cases, it is similar to that of the naive method.
+Since there are settings in simulation studies where the optimal matrix is close to being degenerate, those cases in which the accuracy deteriorates to that of the naïve method may be attributed to the fact that the estimation of $\theta$ is not stable.
+Regardless of whether (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate the ATT without bias.
+Table 1: Comparison of estimates of the ATT when the propensity scores are given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix.
+Est indicates the average of the estimates obtained by each simulation, and Conf Int indicates its empirical 95% confidence interval.
 CBD-id CBD-opt MLE
 $\beta^{*}$ $\alpha^{*}$ $n$ True Est Conf Int Est Conf Int Est Conf Int
-200 0.09 0.09 [--0.12, 0.30] 0.17 [--0.20, 0.51] --0.00 [--0.22, 0.21]
-1.0 400 0.08 0.09 [--0.07, 0.24] 0.17 [--0.14, 0.45] --0.01 [--0.17, 0.14]
-600 0.08 0.08 [--0.03, 0.21] 0.19 [--0.13, 0.12] --0.01 [--0.13, 0.12]
-200 0.10 0.09 [--0.13, 0.32] 0.13 [--0.10, 0.37] --0.14 [--0.43, 0.15]
-0.1 2.0 400 0.10 0.10 [--0.08, 0.25] 0.13 [--0.04, 0.29] --0.13 [--0.34, 0.06]
-600 0.09 0.10 [--0.04, 0.24] 0.13 [--0.01, 0.27] --0.12 [--0.30, 0.03]
-200 0.09 0.11 [--0.15, 0.39] 0.12 [--0.14, 0.37] --0.31 [--0.80, 0.10]
-3.0 400 0.09 0.11 [--0.07, 0.29] 0.13 [--0.06, 0.31] --0.30 [--0.60, --0.03]
-600 0.10 0.11 [--0.03, 0.26] 0.12 [--0.02, 0.27] --0.30 [--0.55, --0.08]
-200 0.42 0.43 [0.21, 0.65] 0.52 [0.09, 0.93] 0.31 [0.06, 0.57]
-1.0 400 0.43 0.43 [0.27, 0.59] 0.53 [0.17, 0.85] 0.31 [0.15, 0.49]
+200 0.09 0.09 [-0.12, 0.30] 0.17 [-0.20, 0.51] -0.00 [-0.22, 0.21]
+1.0 400 0.08 0.09 [-0.07, 0.24] 0.17 [-0.14, 0.45] -0.01 [-0.17, 0.14]
+600 0.08 0.08 [-0.03, 0.21] 0.19 [-0.13, 0.12] -0.01 [-0.13, 0.12]
+200 0.10 0.09 [-0.13, 0.32] 0.13 [-0.10, 0.37] -0.14 [-0.43, 0.15]
+0.1 2.0 400 0.10 0.10 [-0.08, 0.25] 0.13 [-0.04, 0.29] -0.13 [-0.34, 0.06]
+600 0.09 0.10 [-0.04, 0.24] 0.13 [-0.01, 0.27] -0.12 [-0.30, 0.03]
+200 0.09 0.11 [-0.15, 0.39] 0.12 [-0.14, 0.37] -0.31 [-0.80, 0.10]
+3.0 400 0.09 0.11 [-0.07, 0.29] 0.13 [-0.06, 0.31] -0.30 [-0.60, -0.03]
+600 0.10 0.11 [-0.03, 0.26] 0.12 [-0.02, 0.27] -0.30 [-0.55, -0.08]
+1.0 200 0.42 0.43 [0.21, 0.65] 0.52 [0.09, 0.93] 0.31 [0.06, 0.57]
 600 0.43 0.43 [0.31, 0.56] 0.55 [0.20, 0.80] 0.31 [0.18, 0.45]
-200 0.47 0.46 [0.23, 0.70] 0.50 [0.27, 0.75] 0.21 [--0.10, 0.51]
-0.5 2.0 400 0.46 0.47 [0.29, 0.63] 0.50 [0.32, 0.67] 0.21 [--0.00, 0.41]
+0.5 2.0 200 0.47 0.46 [0.23, 0.70] 0.50 [0.27, 0.75] 0.21 [-0.10, 0.51]
 600 0.46 0.47 [0.33, 0.61] 0.51 [0.37, 0.64] 0.22 [0.04, 0.39]
-200 0.47 0.50 [0.24, 0.78] 0.51 [0.25, 0.77] 0.05 [--0.45, 0.47]
-3.0 400 0.47 0.50 [0.32, 0.69] 0.52 [0.33, 0.70] 0.06 [0.32, 0.69]
-600 0.48 0.50 [0.35, 0.66] 0.51 [0.37, 0.66] 0.05 [--0.20, 0.28]
-200 0.85 0.85 [0.62, 1.10] 0.97 [0.47, 1.44] 0.71 [0.44, 1.01]
-1.0 400 0.85 0.85 [0.68, 1.03] 0.98 [0.53, 1.35] 0.70 [0.52, 0.91]
+3.0 200 0.47 0.50 [0.24, 0.78] 0.51 [0.25, 0.77] 0.05 [-0.45, 0.47]
+600 0.48 0.50 [0.35, 0.66] 0.51 [0.37, 0.66] 0.05 [-0.20, 0.28]
+1.0 200 0.85 0.85 [0.62, 1.10] 0.97 [0.47, 1.44] 0.71 [0.44, 1.01]
 600 0.85 0.85 [0.72, 0.99] 1.00 [0.57, 1.31] 0.71 [0.55, 0.87]
-200 0.92 0.93 [0.68, 1.19] 0.97 [0.73, 1.24] 0.63 [0.30, 0.96]
-1.0 2.0 400 0.92 0.93 [0.74, 1.10] 0.97 [0.78, 1.15] 0.64 [0.40, 0.85]
+1.0 2.0 200 0.92 0.93 [0.68, 1.19] 0.97 [0.73, 1.24] 0.63 [0.30, 0.96]
 600 0.92 0.93 [0.78, 1.08] 0.98 [0.83, 1.12] 0.65 [0.46, 0.83]
-200 0.95 0.99 [0.71, 1.28] 1.00 [0.73, 1.27] 0.50 [--0.02, 0.93]
-3.0 400 0.95 0.99 [0.80, 1.18] 1.00 [0.81, 1.20] 0.50 [0.18, 0.80]
+3.0 200 0.95 0.99 [0.71, 1.28] 1.00 [0.73, 1.27] 0.50 [-0.02, 0.93]
 600 0.95 0.99 [0.83, 1.15] 1.00 [0.85, 1.16] 0.50 [0.24, 0.74]
@@
 600 2.85 2.93 [2.71, 3.15] 2.95 [2.75, 3.16] 2.28 [1.97, 2.59]
-4 Model selection criteria for the semiparametric difference-in-difference method when propensity score is known
-4.1 Risk function and its asymptotic evaluation in the semiparametric difference-in-difference method
-As mentioned in the Introduction, there are no reasonable model selection criteria even for the basic setting as treated in Abadie (2005), we first derive model selection criteria when the propensity score is known, and then develop model selection criteria when using the proposed CBD method.
-Specifically, using the ideas of Baba et al.
-(2017) and Baba and Ninomiya (2021), we derive model selection criteria as asymptotically unbiased estimators of the natural risk function corresponding to the loss function used to obtain SDID estimators.
-By using these criteria, the dimension of $\bm{\theta}$ is determined and conponent of $\bm{x}$ is selected.
-As its natural risk function, based on (4), we define a weighted risk function,
+4 Model selection criterion for semiparametric difference-in-differences when propensity scores are known
+4.1 Risk function and its asymptotic evaluation in semiparametric difference-in-differences
+As mentioned in Section 1, there are no reasonable model selection criteria even for the basic setting treated in Abadie (2005).
+To remedy this situation, in this section, we will first derive a model selection criterion for when the propensity scores are known and then develop a model selection criterion for when the proposed CBD method is used.
+Specifically, by using the ideas of Baba et al.
+(2017) and Baba and Ninomiya (2025), we will derive the model selection criteria as asymptotically unbiased estimators of the natural risk corresponding to the loss function used to obtain SDID estimators.
+Using these criteria, the dimension of $\bm{\theta}$ is determined and the component of $\bm{x}$ is selected.
+A natural choice is to define a weighted risk based on (4) as
 $\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
@@
 […]
-According to the derivation of the conventional $C_{p}$ criterion, we remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term.
-For simplicity, we refer to the third term as the bias when the risk is simply estimated.
-If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), then the third term can be written as
+Following the derivation of the conventional $C_{p}$ criterion, we will remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term.
+For simplicity, we will refer to the third term as a bias when the risk is simply estimated.
+If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), the third term can be written as
 $2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})\}\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
-$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
-$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (17)
-The second equality follows from ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$ holds from (3).
-Here, we use the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that used to construct $\hat{\bm{\theta}}$.
-Then, letting $c^{\rm limit}$ be the weak limit of the third term with expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (16).
-Note that we consider a risk function based on the loss function in deriving the SDID estimator, but it is also possible to consider a simple squared error.
+$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (16)
+The second equality is satisfied by ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since it follows from (3) that ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$.
+Here, we have used the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that was used to construct $\hat{\bm{\theta}}$.
+Then, letting $c^{\rm limit}$ be the weak limit of the third term with the expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (15).
+Note that while we consider a risk based on the loss function in deriving the SDID estimator, it is also possible to consider a simple squared error.
 4.2 Derivation of model selection criterion
-In the case of known propensity scores, we denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})=e^{[1]}(\bm{x}_{i})$, $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})=\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity.
-The main term of the error of the SDID estimator is given by
+In the case of known propensity scores, we will denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})$ as $e^{[1]}(\bm{x}_{i})$ and $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})$ as $\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity.
+The main term of the error of the SDID estimator can be expressed as
 $\hat{\bm{\theta}}-\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}$
 $={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\varepsilon_{i}+a(\bm{x}_{i})-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}.$
-Thus, (17) evaluates to
+Accordingly, (16) is evaluated as
 ${\rm E}[c^{\rm limit}]=2\sum_{i=1}^{n}{\rm E}\bigg[\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j})\bm{x}_{j}\{\varepsilon_{j}+a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}\}\bigg].$
-This expectation for the case in which $i\neq j$ is the product of the expectations for $i$ and $j$ from the independence among samples. The expected value for $i$ can be calculated as
+This expectation for the case of $i\neq j$ is the product of the expectations for $i$ and $j$, from the independence among the samples. The expectation for $i$ can be calculated as
 ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]={\rm E}_{\bm{x}_{i}}[{\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]=\bm{0}.$ (17)
-Therefore, we have only to consider the case $i=j$.
-Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from 18, and the following theorem is obtained.
+Therefore, we only have to consider the case of $i=j$.
+Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from (17), and the following theorem is obtained.
 Theorem 2.
-When the propensity score is known,
-${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\}$
-is obtained, where
+When the propensity scores are known, it holds that
+${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]$
+$=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\},$
+where
 $\sigma^{2}(\bm{x})\equiv{\rm E}[\varepsilon^{2}\mid\bm{x}]$ $={\rm E}[\{\rho(\bm{t},\bm{x})\Delta\}^{2}\mid\bm{x}]-{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}.$ (18)
-While empirical evaluation of ${\rm E}[c^{\rm limit}]$ may seem easily possible, addressing the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$ must be considered.
-The reason is that this cannot be empirically evaluated as $1/n\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$.
-In this case, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ for ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$.
-As a result, when the propensity score is known, we propose
+Although one may think that an empirical evaluation of ${\rm E}[c^{\rm limit}]$ is easy, we must consider how to deal with the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$.
+This cannot be empirically evaluated as $n^{-1}\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$.
+Here, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ into ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$.
+As a result, when the propensity scores are known, we propose
 $\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}$
-$+2{\rm tr}\biggl{(}\biggl{\{}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{\}}^{-1}\biggl{[}\sum_{i=1}^{n}\Bigl{\{}\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\Bigr{\}}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{]}\biggr{)}$ (20)
-as the model selection criterion for the SDID method.
-4.3 Extension of the QICw
-In order to provide a comparator for the proposed model selection criterion, we extend the QICw (Platt et al.
-2013) for the setting of this article, which is the model selection criterion for propensity score weighted estimation on missing data.
-First, for the term of goodness of fit, we construct it using the same ideas as for the QICw.
-Specifically, the log-likelihood function with no missing data is differentiated, a weighted potential outcome variable is substituted for the one whose expected value is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted for the parameter.
+$+2{\rm tr}\biggl(\biggl\{\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr\}^{-1}\biggl[\sum_{i=1}^{n}\{\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr]\biggr)$ (19)
+as the model selection criterion for the basic SDID approach.
+4.3 Extension of QICw
+As a comparator for the proposed model selection criterion, we will extend QICw (Platt et al.
+2013), which is an intuitive model selection criterion for weighted estimation using propensity scores, to the setting of this paper.
+First, we construct the goodness-of-fit term using the same idea as for QICw.
+To do so, the log-likelihood function for full data is differentiated, a weighted potential outcome variable is substituted into the one whose expectation is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted into the parameter.
 The number of parameters is used as the penalty term.
-In this setting, the log-likelihood function with no missing data should be
-$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2},$
-and if we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, the outcome variable and the conditional expectation will be the same from (3).
-Then, the sum of the goodness of fit and penalty terms multiplied by the constant $2\sigma^{2}$, that is,
+Within this setting, the log-likelihood function for full data is
+$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2}.$
+If we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, its conditional expectation can be shown to be the same as the conditional expectation of the outcome variable from (3).
+We will call the sum of the goodness-of-fit term and the penalty term multiplied by a constant $2\sigma^{2}$, that is,
 $\sum_{i=1}^{n}\{\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}+2\sigma^{2}p,$
-we call QICw for SDID.
-In fact, if $\sigma^{2}(\bm{x})$ is constant in our proposed criterion, then it is equivalent to the QICw.
-The problem is to estimate the variance $\sigma^{2}$ of “$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$”.
-This estimated value cannot be obtained from the data using only the present assumptions.
-Since $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ holds from $y^{[0]}(0)=y^{[1]}(0)$, and from Assumption 1, the expected value of “$\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$” is the same as the expected value of “$\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$”.
-Since the variance must be evaluated, we further assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, respectively, and calculate that
+the QICw for the SDID approach.
+In fact, if $\sigma^{2}(\bm{x})$ is a constant, then our proposed criterion is equivalent to the QICw.
+The problem is to estimate the variance $\sigma^{2}$ of the difference, $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$.
+This estimate cannot be obtained from the data using only the present assumptions.
+Since it follows from $y^{[0]}(0)=y^{[1]}(0)$ that $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$, and from Assumption 1, the expectation of $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$ is the same as the expectation of $\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$.
+However, in order to evaluate the variance, we will assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, from which we have
 $\sigma^{2}={\rm Var}[\Delta^{[1]}\mid d^{[1]}=1]+{\rm Var}[\Delta^{[0]}\mid d^{[0]}=1]$
-$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}$
-From the above, we define $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, and
+$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}.$
+Moreover, defining $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, we will use
 $\hat{\sigma}^{2}=\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}^{2}-\bigg(\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}\bigg)^{2}+\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}^{2}-\bigg(\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}\bigg)^{2}$
-is used in the present study.
+in the present study.
 Remark 3.
-Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (19), if all expected values do not actually depend on $\bm{x}$, then
-$\sigma^{2}(\bm{x})=$ ${\rm E}\bigg{[}\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg{|}\ d^{[1]}=1\bigg{]}-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg{[}\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg{|}\ d^{[0]}=1\bigg{]}-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
+Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (18), if all expectations do not actually depend on $\bm{x}$, then
+$\sigma^{2}(\bm{x})=\$${\rm E}\bigg[\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg|\ d^{[1]}=1\bigg]-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg[\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg|\ d^{[0]}=1\bigg]-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
 $+2{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]{\rm E}[\Delta^{[0]}\mid d^{[0]}=1].$
@@
 4.4 Numerical experiments
-By simulations, we evaluate the performance of the proposed criterion (20) when the propensity score is known. As data generation settings, we consider the Case 1-1,
+Here, the performance of the proposed criterion (19) when the propensity scores are known is evaluated in simulation studies. The data generation settings include Case 1-1,
 $x_{1}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
@@
 $y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]},$
-and the Case 1-2,
+and Case 1-2,
 $x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
 $y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
-$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]}.$
-In both cases, $\beta^{*}$ represents the degree of contribution of the covariate $x_{1}$ to the causal effect.
-In Table 2, for some $\beta^{*}$ and sample size $n$, we examine whether the second term of (20), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate.
-In all cases, our penalty terms are close to the true values of the bias, while the penalty terms in the QICw significantly underestimate the bias.
-Table 2: Bias evaluation when the propensity score is known: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method.
-The penalty term of the QICw is also given as a reference.
+$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}(x_{1}+x_{2})+\epsilon^{[1]}.$
+In both cases, $\beta^{*}$ represents the magnitude of contribution of the covariate $x_{1}$ to the causal effect.
+Table 2 shows whether the second term in (19), the penalty term of the proposed criterion, accurately approximates the bias in (16) that we originally wanted to evaluate.
+For all values of $\beta^{*}$ and sample size $n$, our penalty terms are close to the true values of the bias, while the evaluated penalty terms in the QICw significantly underestimate the bias.
+Table 2: Bias evaluation for the case where the propensity scores are known: True denotes the true value of the bias evaluated by the Monte Carlo method, and Proposal the evaluated penalty term of the proposed criterion.
+The evaluated penalty term of the QICw is listed as a reference.
 Case 1-1 Case 1-2
@@
 600 367.73 365.75 5.22 435.32 430.09 23.87
-Next, we examine the performance of model selection using the proposed criteria. Data is generated from
+Next, let us examine the performance of the model selection using the proposed criterion, with data generated from
 $x_{1},x_{2},\ldots,x_{l}\sim{\rm Uniform}(0,2),\quad\bm{x}=(1,x_{1},x_{2},\ldots,x_{l})^{{\rm T}},$
 $y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
-$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{\beta}^{[1]*}{}^{{\rm T}}\bm{x}+\epsilon^{[1]},$
-and we consider the Case 2-1,
+$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{x}^{{\rm T}}\bm{\beta}^{[1]*}+\epsilon^{[1]},$
+in Case 2-1,
 $l=4,\quad\bm{\beta}^{[1]*}=(1,\beta^{*},\bm{0}_{3}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
-the Case 2-2,
+Case 2-2,
 $l=4,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{2}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
-and the Case 2-3,
+and Case 2-3,
 $l=6,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{4}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})).$
-Table 3 compares the proposed criterion and the QICw by selecting models using the forward selection method.
-As the main index to measure the goodness of the criteria, we use the empirical estimator of the risk function, specifically, the average of 3000 calculations of
+Table 3 compares the proposed criterion and the QICw when using the forward selection method.
+As the main index to measure the goodness of the criteria, we use an empirical estimate of the risk, specifically, the average of 3,000 values of
 $\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha})(\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}.$
-In the Case 2-2 setting, where there are relatively few unnecessary covariates among the candidates, that is, the negative effect of false positives is less likely to appear, the QICw slightly exceeds the proposed criterion in some cases for $\beta^{*}=0.5$ and $\beta^{*}=1.0$.
-Otherwise, the proposed criterion outperforms the QICw, especially in settings with many covariates unrelated to the true structure such as Case 2-1 and Case 2-3, a clear difference is observed.
-Since the QICw underestimates the penalty term, the false positive value must be large and too many extraneous variables must be selected in all settings.
-Table 3: Comparison of the proposed criterion and the QICw through model selection when the propensity score is known: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
+In Case 2-2, where there are relatively few unnecessary covariates among the candidates, that is, where the negative effect of false positives is less likely to appear, the QICw slightly outperforms the proposed criterion for $\beta^{*}=0.5$ and $\beta^{*}=1.0$.
+However, the proposed criterion outperforms the QICw for the other values of beta in Case 2-2, and it outperforms the QICw for all values in Case 2-1 and Case 2-3, i.e., settings with many covariates are unrelated to the true structure.
+Here, since the QICw underestimates the penalty term, the false positive must be large and too many covariates are selected in all settings.
+Table 3: Comparison of the proposed criterion and the QICw through model selection for case in which the propensity scores are known: Risk denotes the empirical estimate of the risk of the selected model, TP the average number of covariates that are true positive, and FP the average number of covariates that are false positive.
 Case 2-1 Case 2-2 Case 2-3
@@
 QICw 67.51 1.00 1.91 135.24 2.00 1.17 191.42 2.00 2.36
-5 Model Selection Criteria for the CBD Method
+5 Model selection criterion for covariate balancing difference-in-differences
 5.1 Derivation of model selection criterion
-As in the case with known propensity scores in Section 4.2, we derive the model selection criteria when using the CBD method. We evaluate
+As we did for the case of known propensity scores in Section 4.2, here, we will derive a validated model selection criterion for when the CBD method is used. Let us evaluate
 $\sum_{i=1}^{n}2{\rm E}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-{\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]\}\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*})],$
-which is the third term of the risk function (17).
-First, substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ for ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$, and then substituting the right-hand side of (15) for $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$, this third term can be written as
-$2{\rm E}\bigg{[}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}$
+which is the third term of the risk in (16).
+Substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ into ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$ first and then substituting the right-hand side of (13) into $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$ yields
+$2{\rm E}\bigg[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}\bigg\{\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}\bigg\}^{-1}$
 […]
-Since $(1/n)\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges to
-$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}\{e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}\}$
-in probability, if we let $\bm{u}$ be the normal random vector that is the weak limit of
+Since $n^{-1}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges in probability to
+$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}],$ (20)
+by letting $\bm{u}$ be the Gaussian random vector that is the weak limit of
 $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i},$ (21)
-then the weak limit of (17) removing the expectation is expressed as
+we can express the weak limit of (16) with expectation removed as
 $c^{\rm limit}=2\bm{u}^{{\rm T}}\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}\bm{u}^{{\rm T}}\}.$
-In the setting where the estimator $\hat{\bm{\theta}}^{\rm CBD}$ is consistent with $\bm{\theta}^{*}$, ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ holds because the expected value of $\bm{u}$ becomes $\bm{0}$, and the asymptotic evaluation of (17), which is to be used in the information criterion, is expressed as
+When $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator of $\bm{\theta}^{*}$, it holds that ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ since the expectation of $\bm{u}$ becomes $\bm{0}$.
+Accordingly, the asymptotic evaluation of (16), which is to be used in the information criterion, can be written as
 ${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}{\rm Var}[\bm{u}]\}.$
-Here, we write down ${\rm Var}[\bm{u}]$. As the error of the parameter $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity score can be shown to be (14), using this, (21) can be expanded as
+Now let us derive a specific expression for ${\rm Var}[\bm{u}]$. Since the error of the estimator $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity scores is written as (12), we can expand (21) as
 $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
 $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bm{x}_{i}\bigg\{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
-$\ \ \ \ \ +\dfrac{\partial}{\partial\bm{\alpha}^{{\rm T}}}\bigg{(}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bigg{[}\frac{d^{[1]}_{i}-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{]}\bigg{)}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})\bigg{\}}\{1+{\rm o}_{\rm P}(1)\}$
-$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
-$\ \ \ \ \ \ \ \ \ \ -\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Big{]}\{1+{\rm o}_{\rm P}(1)\},$
+[…]
+[…]
+$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
+[…]
 where
 $\bm{M}(\bm{\alpha},\bm{\theta})={\rm E}\biggl[\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x};\bm{\alpha})\biggl[\frac{d^{[1]}-1}{\{1-e^{[1]}(\bm{x};\bm{\alpha})\}^{2}}\Delta-\bm{x}^{{\rm T}}\bm{\theta}\biggr]\bm{x}\biggr].$
-From the above, by defining
+Then, by defining
 $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\equiv{\rm Var}$ $[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}$
 $-\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})],$ (22)
-it can be shown to be ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
+it can be shown that ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
 Theorem 3.
-When the propensity score is estimated by the CBD method, the third term of (16) is evaluated as
-${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}.$
-As a result, when the propensity score is estimated by the CBD method, we propose
-$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})(\rho(\bm{d}_{i},\bm{x}_{i})(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD})^{2}+2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}$ (23)
-as the model selection criterion for the SDID method, which is an asymptotically unbiased estimator of (16).
-Also in this case, the penalty term depends on the true values of the parameters, thus empirical estimates are used in practice.
-Specifically, $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ is used instead of $\bm{L}(\bm{\alpha}^{\dagger})$ and
-$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\equiv\frac{1}{n}$ $\sum_{i=1}^{n}\Bigl{[}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
-$-\bm{M}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Bigr{]}^{2}$
-is used instead of $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$, where
-$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\biggl{(}\frac{d^{[1]}_{i}-1}{(1-e^{[1]}(\bm{x}_{i};\bm{\alpha}))^{2}}\Delta_{i}-\bm{x_{i}}^{{\rm T}}\bm{\theta}\biggr{)}\bm{x_{i}}.$
-Here, the propensity score is estimated using the covariate balancing approach, but the information criterion can be derived in the same way when the MLE is used (see Supplementary material).
+When the propensity scores are estimated using the CBD method, the third term in (15) is evaluated as
+${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\},$
+where $\bm{L}(\cdot)$ and $\bm{V}(\cdot,\cdot)$ are defined in (20) and (22), respectively.
+Since the penalty term in this theorem depends on the true values of the parameters, we can use empirical estimates in practice.
+Specifically, we will replace $\bm{L}(\bm{\alpha}^{\dagger})$ by $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ and $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$ by
+$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})$
+$\equiv\frac{1}{n}\sum_{i=1}^{n}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
+[…]
+where
+$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bigg[\frac{d^{[1]}_{i}-1}{\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha})\}^{2}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}\bigg]\bm{x}_{i}.$ (23)
+As a result, for when the propensity scores are estimated with the CBD method, we propose
+$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}^{2}+2{\rm tr}\{\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})^{-1}\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\}$ (24)
+as the model selection criterion, which is an asymptotically unbiased estimator of the risk in (15).
+Here, the propensity scores are estimated using covariate balancing, but the information criterion can be derived in the same way when the MLE is used (see Supplementary Material).
 5.2 Numerical experiments
-Through simulations, we examine the accuracy of the approximation of the penalty term and the performance of the model selection criterion (23) based on Theorem 3 for the proposed CBD method.
-As the weighting matrix in the generalized method of moments, we use the identity matrix based on the results of Section 4.4.
-But even when an optimal matrix is used as the weighting matrix, the accuracy of the approximation of the penalty term and the performance of the model selection are comparable to those obtained when the identity matrix is used (see Supplementary Material).
-In Table 4, data are generated as in Case 1-1 and Case 1-2 of Section 4.4 to verify whether the second term of (23), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate for some parameter values $\beta^{*}$ and sample size $n$.
-In all cases, the penalty term of the proposed criterion provides values much closer to the bias than the penalty term of the QICw, and its approximation accuracy is confirmed to be high.
-Conversely, in all cases, the penalty term of the QICw significantly underestimates the bias.
-Table 4: Bias evaluation when the propensity score is estimated by the CBD method with identity matrix as the weighting matrix: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method.
-The penalty term of the QICw is also given as a reference.
+The simulation studies presented in this section examine the accuracy of the approximation for the penalty term and the performance of the model selection criterion in (24) for the proposed CBD method.
+On the basis of the results in Section 4.4, we will use the identity matrix as the weighting matrix in the generalized method of moments and note that, even when the optimal matrix is used as the weighting matrix, the accuracy of the approximation for the penalty term and the performance of the model selection with comparable to those obtained with the identity matrix (see the Supplementary Material).
+Table 4 lists results for data generated as in Cases 1-1 and Cases 1-2 in Section 4.4 to check whether the second term in (24), the penalty term of the proposed criterion, accurately approximates the bias in (16), which we originally wanted to evaluate, for various values of the parameter $\beta^{*}$ and sample size $n$.
+In all cases, the evaluated penalty terms of the proposed criterion are much closer to the bias than the evaluated penalty terms of the QICw, and its approximation accuracy is high.
+In contrast, the penalty term of the QICw significantly underestimates the bias in all cases.
+Table 4: Bias evaluation for case in which the propensity scores are estimated using the CBD method with the identity matrix as the weighting matrix: True denotes the true value of the bias evaluated by the Monte Carlo method, and Proposal the penalty term of the proposed criterion.
+The penalty term of the QICw is also listed for reference.
 Case 1-1 Case 1-2
@@
 600 17.88 18.49 5.22 53.55 54.64 23.90
-In Table 5, data are generated as in Case 2-1, Case 2-2 and Case 2-3 of Section 4.4 and the proposed criterion and the QICw are compared by model selection using the forward selection method.
-The main index of the goodness of the model selection criteria is again the empirical estimate of the risk.
-Unlike the results in Section 4.4, where the propensity score is known, the proposed criterion always outperforms the QICw, even in Case 2-2, where relatively few covariates are unnecessary among the candidates.
-On the other hand, as in Case 2-1 and Case 2-3, the proposed criterion is particularly superior when the candidates include many covariates that do not contribute to the true structure, which is similar to Section 4.4.
-The penalty term is underestimated in the QICw, therefore, too many extraneous variables must be selected.
-Table 5: Comparison of the proposed criterion and the QICw through model selection when the propensity score is estimated using the CBD method with identity matrix as the weighting matrix: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
+Table 5 lists results for data generated as in Case 2-1, Case 2-2, and Case 2-3 in Section 4.4 and comparing the proposed criterion and the QICw through the forward selection method.
+The main index of goodness for the model selection criteria is again the empirical estimate of the risk.
+Unlike Section 4.4, where the propensity scores are known, the proposed criterion always outperforms the QICw in all cases, even in Case 2-2, where relatively few covariates are unnecessary among the candidates.
+The proposed criterion is especially superior when the candidates include many covariates that do not contribute to the true structure, as in Case 2-1 and Case 2-3 of Section 4.4.
+The QICw underestimates the penalty term; thus, too many covariates are selected.
+Table 5: Comparison of the proposed criterion and the QICw through model selection for case in which the propensity scores are estimated using the CBD method with the identity matrix as the weighting matrix: Risk denotes the empirical estimate of the risk of the selected model, TP the average number of covariates that are true positive, and FP the average number of covariates that are false positive.
 Case 2-1 Case 2-2 Case 2-3
@@
 6 Real data analysis
-The LaLonde dataset is treated in LaLonde (1986) and included in the R package Matching version 4.10.14 (Sekhon 2011).
-The group that took the U.S.
+The LaLonde dataset, originally presented in LaLonde (1986), is included in the R package Matching version 4.10.14.
+In the dataset, the group that took the U.S.
 job training program in 1976 is denoted as $t=1$ and the group that did not take the program is denoted as $t=0$.
 The difference in annual income in 1978 after the training for each group is estimated via the ATT.
-Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are treated as covariates, i.e.
+Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are covariates, i.e.
 $p=7$, and the outcome variable is the difference between income in 1978 (re78) and income in 1974 (re78).
 The sample size is $n=445$.
-The objective of this section is not to extract information from this data, but to examine whether the proposed method makes a significant difference from existing methods, and therefore, a sufficiently small number of samples are assumed.
-Specifically, for a natural number $k$, the $23k-$ 2 rows of data are analyzed in Block 1, $13k-$ 1 rows in Block 2, and $3k$ rows in Block 3, respectively.
-Regarding the estimation, we first estimate the propensity score using the CBD method with the identity matrix as the weighting matrix.
-Then, assuming that the ATT can be written as a linear sum of covariates, we estimate its coefficients by the SDID method.
-Table 5 selects covariates using the proposed criterion or the QICw with the forward selection method and estimates the coefficients of the selected variables.
-The QICw selects all variables in any Block, while the proposed criterion does not select some variables.
-In particular, the difference between the models selected for Block 1 and Block 3 is remarkable.
-Because we do not know the true structure of the real data, it is impossible to judge the superiority or inferiority of the two criteria, however, we can see that the difference between them is quite large.
-In a setting such as this article, it is essential to use model selection criteria that are theoretically valid rather than intuitive ones.
-Table 6: Comparison of the proposed criterion and the QICw when propensity scores are estimated using the CBD method with the identity matrix as the weighing matrix, through model selection experiments on LaLonde data.
-The numbers with exponential notation are the estimated coefficients for each variable, and a 00 means that the variable was not selected in the model selection.
+The objective of our experiment here is not to extract information from these data, but to determine whether the proposed method makes a significant difference from an existing method.
+Therefore, to reduce the sample size of the dataset, we divided the data into three parts and analyzed each of them.
+Specifically, for every integer $k$, the 3k23k-2-th row of data are included in Block 1, the 3k13k-1-th row in Block 2, and the $3k$-th row in Block 3, respectively.
+First, we estimate the propensity scores by using the covariate balancing with the identity matrix as the weighting matrix.
+Then, supposing that the ATT can be written as a linear sum of covariates, we estimate their coefficients by using the SDID approach.
+In Table 5, the covariates are selected using the proposed criterion or the QICw with the forward selection method and the coefficients of the selected covariates are estimated.
+The QICw selects all covariates in any block, while the proposed criterion does not select some variables.
+In particular, the difference between the models selected for Block 1 and Block 3 is rather large.
+Because we do not know the true structure of the real data, it is impossible to judge which of the two criteria is best; however, we can see that the difference between them is quite large.
+These results indicates that, in a setting such as here, it is essential to use a model selection criterion that is theoretically valid rather than an intuitive one.
+Table 6: Comparison of the proposed criterion and the QICw through model selection for the LaLonde dataset.
+The numbers with exponential notation are the estimated coefficients for each covariate, and “0” means that the covariate is not selected in the model selection.
 Block intercept age educ re74
-1 Proposal --2.14E+3 0E+0 0E+0 0E+0
-QICw --1.37E+4 5.29E+2 4.43E+1 --2.01E-1
+1 Proposal -2.14E+3 0E+0 0E+0 0E+0
+QICw -1.37E+4 5.29E+2 4.43E+1 -2.01E-1
 2 Proposal 1.59E+4 0E+0 1.08E+3 2.54E-1
-QICw 2.38E+4 --1.91E+2 8.22E+2 2.73E-1
-3 Proposal --8.50E+2 0E+0 0E+0 0E+0
-QICw --8.51E+3 6.13E+0 3.35E+2 2.56E-1
+QICw 2.38E+4 -1.91E+2 8.22E+2 2.73E-1
+3 Proposal -8.50E+2 0E+0 0E+0 0E+0
+QICw -8.51E+3 6.13E+0 3.35E+2 2.56E-1
 Block black hisp married nodegr
 1 Proposal 0E+0 0E+0 0E+0 0E+0
-QICw 3.55E+2 --8.19E+3 --9.41E+2 --1.30E+3
-2 Proposal --1.86E+4 --2.34E+4 4.02E+3 --1.26E+4
-QICw --1.85E+4 --2.42E+4 4.88E+3 --1.34E+4
-3 Proposal 0E+0 --1.03E+4 0E+0 0E+0
-QICw 3.84E+3 --6.71E+3 2.92E+3 --5.90E+2
+QICw 3.55E+2 -8.19E+3 -9.41E+2 -1.30E+3
+2 Proposal -1.86E+4 -2.34E+4 4.02E+3 -1.26E+4
+QICw -1.85E+4 -2.42E+4 4.88E+3 -1.34E+4
+3 Proposal 0E+0 -1.03E+4 0E+0 0E+0
+QICw 3.84E+3 -6.71E+3 2.92E+3 -5.90E+2
 7 Discussion
-In this article, we have proposed the CBD method, a SDID method incorporating covariate balancing.
-In order to ensure the doubly robustness, i.e., that the ATT can be consistently estimated if the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly assumed, we have shown that it is sufficient to balance the distribution up to the second-order moment of the covariates across groups with the treatment group as the target population.
-This fact is important to note because covariate balancing usually balances first-order moment.
-We also have developed model selection criteria for the CBD method as an asymptotically unbiased estimator of the risk based on the loss function used for the estimation.
-Specifically, we have derived model selection criteria using a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference, and we have shown that the penalty term is considerably different from twice the number of parameters that often appears in AIC-type model selection criteria.
-Since there have been no model selection criteria for even the basic SDID method, we have developed the model selection criteria not only for the CBD method, but also for when the propensity score is known or the propensity score is estimated by the MLE.
-Numerical experiments showed that the proposed methods estimate ATT robustly compared to the method that estimates propensity scores by the MLE, and that the proposed criteria clearly reduce the risk targeted by the SDID method compared to the intuitive generalization of the existing information criterion to the SDID method.
-In addition, the real data analysis has confirmed that there were large differences between the results of the proposed method and the existing methods.
-Although this article deals with a linear model of covariates as a simple working model for the outcome, it should be noted that the results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method.
-In fact, Hazlett (2020) treats the function $\phi(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in the reproducing kernel Hilbert space as $q$ may be larger than $p$ or $n$.
-And they assume that the outcome is represented by a linear model of $\phi(\bm{x})$, the feature map of $\bm{x}$, which in our setting can be written as
-${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\theta}.$ (24)
+Our proposal, the CBD method, is an SDID approach incorporating covariate balancing.
+In order to ensure double robustness, i.e.
+to ensure that the ATT is consistently estimated when the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly supposed, we have shown that the second-order moments of the distribution for the covariates should be balanced between the treatment and control groups with the former as the target population.
+Since covariate balancing usually balances the first-order moments, the value for the fact that the moment is second-order rather than first-order lies more in having been discovered due to its unexpectedness than in how to demonstrate its validity.
+Moreover, our model selection criterion for the CBD method is an asymptotically unbiased estimator of risk based on the loss function used for the estimation.
+Specifically, we derived a model selection criterion using just a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference.
+The penalty term has been shown to be considerably different from twice the number of parameters that often appears in the AIC-type model selection criteria.
+Since no model selection criteria existed for even the basic SDID approach, we developed a model selection criterion not only for the CBD method, but also for when the propensity scores are known or estimated by MLE.
+Numerical experiments show that the proposed method estimates the ATT more robustly than the method that estimates propensity scores by using MLE and that the proposed criterion clearly reduces the risk targeted by the SDID approach compared with the intuitive generalization of the existing information criterion to the SDID approach.
+In addition, analyses of real data confirm that there is a large difference between the results of the proposed method and those of the existing method.
+Although this paper deals with a simple linear model of covariates as a working model for the outcome, its results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method.
+In fact, Hazlett (2020) has treated the function $\bm{\phi}(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in a reproducing kernel Hilbert space where $q$ may be larger than $p$ or $n$.
+They assumed that the outcome could be represented by a linear model of $\bm{\phi}(\bm{x})$, which in our setting can be written as
+${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\bm{\phi}(\bm{x})^{{\rm T}}\bm{\theta}.$ (25)
 Then, by using
-$\sum_{i=1}^{n}\bigg{\{}\frac{d^{[1]}_{i}\phi(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\phi(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg{\}}=0$
-as the moment condition for estimating the propensity score, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear.
-In the setting of this article, we also consider the situation (24) and when making the estimation doubly robust, we assume that the change in the outcomes follows
-${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
-Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\phi(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semidefinite symmetric matrix whose $(l,m)$ component is $k(x_{l},x_{m})$.
+$\sum_{i=1}^{n}\bigg\{\frac{d^{[1]}_{i}\bm{\phi}(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\bm{\phi}(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg\}=\bm{0}$
+as the moment conditions for estimating the propensity scores, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear.
+For making a doubly robust estimation using the kernel balancing in the setting of this paper, together with the assumption in (25), let us additionally suppose that the change in the outcomes follows
+${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{\phi}(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
+Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\bm{\phi}(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semi-definite symmetric matrix whose $(l,m)$-component is $k(x_{l},x_{m})$.
 If we estimate the parameter $\bm{\alpha}$ that satisfies
@@
 $\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{K}\bigg]=\bm{{\rm O}},$
-and then estimate $\bm{\theta}$ or derive model selection criteria based on it, that would be a natural extension of the proposed method in this article.
-There are other valuable directions for extension.
-For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) propose a DID method in a more flexible setting than the previous designs, where there are multiple time points and the treatment may be initiated at any of them.
-Moreover, De Chaisemartin and D’haultfœuille (2023) develops a DID method when there are more than two groups, and Richardson et al.
-(2023) proposes the generalized DID to estimate causal effects under alternative assumptions to the parallel trend assumption.
-While the outcome in the basic DID method is continuous, Tchetgen et al.
-(2024) deals with the case where the response follows a generalized linear model.
-Thus, the DID method has been theoretically developed in accordance with real problems, and extensions of the proposed method to these problems are recent challenges.
-In econometrics, the synthetic control method (Abadie et al.
-2010) is a major method for causal inference as well as the DID method, and a method combining it with the DID method has also been developed as synthetic DID (Arkhangelsky et al.
+and then estimate $\bm{\theta}$ and derive a model selection criterion based on it, it would be a natural extension of the proposed method.
+There are other valuable directions of study.
+For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) proposed a DID approach in a more flexible setting than the original one, where there are multiple time points and treatment can be initiated at any of them.
+Moreover, De Chaisemartin and D’haultfœuille (2023) developed a DID approach for when there are more than two groups, and Richardson et al.
+(2023) proposed one to estimate causal effects under alternative assumptions to the parallel trend assumption.
+While the outcome in the basic DID approach is continuous, Tchetgen et al.
+(2024) dealt with the case where the outcome follows a generalized linear model.
+In this way, the DID approach has been theoretically developed in accordance with actual problems, so we believe that extending the proposed method to enable it to deal with those problems would thus be a potentially profitable goal for the near future.
+In econometrics, synthetic control (Abadie et al.
+2010) has also become standard in causal inference, and a combination of the two has been developed as synthetic DID (Arkhangelsky et al.
 2021).
-In this method, the objective of the information criterion is what to synthesize and with what weights, and the proposed method cannot necessarily be extended in a similar way, but it is an important topic.
-Furthermore, propensity score analysis with high-dimensional confounding variables is also a hot topic (e.g., Belloni et al.
+Here, because the objective of the information criterion is what to synthesize and with what weights, the proposed method cannot necessarily be used; however, an investigation along the lines presented in this paper might prove fruitful.
+Furthermore, propensity score analysis with high-dimensional covariates is also an important topic (e.g., Belloni et al.
 2017, Chernozhukov et al.
 2018, Ning et al.
-2020b), and even in that analysis, robustly estimating the influence of covariates that directly affect the causal effect and selecting covariates that truly affect the causal effect is an essential task to ensure the validity of the analysis.
-The asymptotic properties of the estimator of causal effects tend to be obtained even when the covariates are high-dimensional as well, and it is surely possible to extend our proposed method for these analyses.
+2020b), wherein selecting covariates that truly have impact on causal effects and robustly estimating these effects is an essential task to ensure the validity of the analysis.
+Even when the covariates are high-dimensional, asymptotic properties of the estimator of causal effects tend to be obtained as well, so it is surely possible to extend our method.
+Disclosure statement
+No potential conflict of interest was reported by the authors.
+Data availability statement
+The LaLonde dataset is available in the R package Matching version 4.10.14 (Sekhon 2011). Example R scripts used for the simulations and real data analyses are available from the authors.
+Funding
+Yoshiyuki Ninomiya gratefully acknowledges support from JSPS KAKENHI (grant numbers 23H00809 and 23K18471)